### v1

# Article Category

Abstract
Deploying autonomous robots in safety-critical applications requires safety guarantees. Provably safe reinforcement learning is an active field of research which aims to provide such guarantees using safeguards. These safeguards should be integrated during training to prevent a large sim-to-real gap. While there are several approaches for safeguarding sampling-based reinforcement learning, analytic gradient-based reinforcement learning often achieves superior performance and sample efficiency. However, there is no safeguarding approach for this learning paradigm yet. Our work addresses this gap by developing the first effective safeguard for analytic gradient-based reinforcement learning. We analyse existing, differentiable safeguards, adapt them through modified mappings and gradient formulations, and integrate them with a state-of-the-art learning algorithm and a differentiable simulation. We evaluate how different safeguards affect policy optimisation using numerical experiments on two classical control tasks. The results demonstrate safeguarded training without compromising performance.
{IEEEkeywords}
Safe reinforcement learning, policy optimisation, differentiable simulation, gradient-based methods, constrained optimisation, first-order analytic gradient-based reinforcement learning
footnotetext: Equal contribution
1 INTRODUCTION
The transfer of physical labour from humans and human-operated machines to robots is a long-standing goal of robotics research. While robots have been successfully deployed in controlled environments, such as factories, deploying strong robots in human proximity is still a challenge [StrongRobot, StrongRobot2]. One reason is a lack of safety guarantees to ensure that robots do not harm humans or themselves [SafetyRequired].
Classical control algorithms, such as PID or MPC controllers, can be designed for safety [SafePID, RobustMPC, MPC2]. While reinforcement learning controllers do not provide safety guarantees, they often achieve state-of-the-art performance [GreatRL, GreatRL2] due to their broad applicability, representation power, and fast inference. In recent years, provably safe reinforcement learning has emerged as a research field to provide safety guarantees [ProvablySafeSurvey, ProvablySafeRL].
Current safeguards are applied in conjunction with reinforcement learning algorithms that rely on the policy gradient theorem to estimate reward landscapes [ActionReplacement1, ActionProjection, RayMasking, BoundaryProjectionLoss, Gauge, BoundaryProjection]. With the advent of differentiable physics simulators [Mujoco, brax, ChainQueen, Thuerey, SAPO], analytically computing the gradient of the reward with respect to the actions has become possible. Reinforcement learning algorithms that exploit these gradients promise faster training and better performance [GradientRL1, GradientRL2, PODS, SHAC]. While these simulators require approximations to remain differentiable, maintaining simulation accuracy is possible. As a result, one can train the policy exclusively in simulation before deploying it to real systems and avoid costly real-world training [Real1, Real2].
Applying safeguards during training is desirable for safety-critical systems to prevent an effectively enlarged sim-to-real gap [simtoreal1, simtoreal2]. Otherwise, the optimisation may converge to strategies that rely on unsafe states or actions. When deployment safeguards subsequently restrict these strategies, the policy can become suboptimal or even fail for non-convex objective landscapes, as it has not learned alternative, safe solutions [DuringTraining]. While crafting reward functions that reliably encode safety requirements could theoretically prevent performance degradation in deployment, this is notoriously difficult without introducing unintended incentives [RewardHacking1, RewardHacking2]. Moreover, safeguards can guide the exploration process to aid learning in challenging solution spaces [RayMasking, PostTraining]. However, there are currently no safeguarding attempts for analytic gradient-based reinforcement learning.
Our work addresses this by combining state-of-the-art analytic gradient-based reinforcement learning algorithms, differentiable safeguards, and differentiable simulations [^1] Code available at https://github.com/TimWalter/SafeGBPO. To ensure safety, we incorporate a range of existing set-based safeguarding methods that map policy actions into a safe action set. We formalise desirable safeguarding properties in the context of differentiable optimisation and analyse existing methods with respect to these criteria. Based on this analysis, we propose targeted modifications, such as custom backward passes or adapted maps, that enhance their suitability for analytic gradient-based reinforcement learning. We also extend the applicability of one of the safeguards to state constraints state sets.
We evaluate the approaches in differentiable simulations of various classical control problems. Throughout, we observe sample efficiency and final performance that exceeds or is on par with unsafe training and sampling-based baselines.
In summary, our core contributions are:
  • the first provably safe analytic policy optimisation,
  • an in-depth analysis of some suitable safeguards,
  • adapted backward passes, an adapted mapping and extended applicability of these safeguards,
  • and an evaluation on two classical control tasks, demonstrating the potential of provably safe reinforcement learning from analytic gradients.
2 RELATED WORK
We provide a literature review on the most relevant research areas: analytic gradient-based reinforcement learning, safeguards, and implicit layers.
2.1 ANALYTIC GRADIENT-BASED REINFORCEMENT LEARNING
Analytic gradient-based reinforcement learning relies on a continuous computational graph from policy actions to rewards, which allows computing the first-order gradient of the reward with respect to the action. Relying on first-order gradient estimators often results in less variance than zeroth-order estimators [GradientOrderDoubt], which are usually obtained using the policy gradient theorem. Less variance leads to faster convergence to local minima of general non-convex smooth objective functions [GradientRL1, GradientRL2]. However, complex environments may lead to optimisation landscapes that are stiff, chaotic or contain discontinuities, which can stifle performance as first-order gradients suffer from empirical bias [GradientOrderDoubt]. Using a smooth surrogate to approximate the underlying noisy reward landscape can alleviate this issue [SHAC]. Moreover, naively backpropagating through time [BPTT] can lead to vanishing or exploding gradients in long trajectories [BPTTProblems].
Various approaches have been introduced to overcome this issue: Policy Optimisation via Differentiable Simulation [PODS] utilises the gradient provided by differentiable simulators in combination with a Hessian approximation to perform policy iteration. It outperforms sampling-based methods. Short-Horizon Actor-Critic (SHAC) [SHAC] tackles the empirical bias of first-order gradient estimators by training a smooth value function through a mean-squared-error loss, with error terms calculated from the sampled short-horizon trajectories through a td-$\lambda$ formulation [RLBible]. It prevents exploding and vanishing gradients by cutting the computational graph deliberately after a fixed number of steps and estimating the terminal value by the critic. The algorithm shows applicability even in contact-rich environments, which tend to lead to stiff dynamics. The successor Adaptive Horizon Actor-Critic [AHAC] has a flexible learning window to avoid stiff dynamics and shows improved performance across the same tasks. Short-Horizon Actor-Critic also inspired Soft Analytic Policy Optimisation [SAPO], which integrates maximum entropy principles to escape local minima.
2.2 SAFEGUARDS
Safeguards are generally categorised according to their safety level [Categorisation, ProvablySafeRL]. Since we seek guarantees, we limit the discussion to hard constraints. Moreover, safeguards must define a differentiable map from unsafe to safe action to allow for backpropagation.
Analytic differentiation and assessing safety necessitate some form of model. However, having the exact model of the system is usually not required. Instead, one can derive conservative approximations of safety based on available data. The two most common techniques to do so are control barrier functions, which guarantee the forward invariance of a set called a safe set [CBF] and reachability analysis, where all possible system behaviours are captured by a reachset-conformant model [Reachability]. Handy safe action sets are robust control invariant sets as they are time-invariant at the cost of potentially conservative results. They can be approximated automatically for nonlinear systems from the state and input constraints of the dynamics [RCI-Sets].
Only maps between continuous action spaces can be differentiable. Krasowski et al. [ProvablySafeRL] present continuous action projection with safe action sets represented by intervals where straightforward re-normalisation is employed to map from the feasible action set. Stolz et al. [RayMasking] generalise this to more expressive sets with their ray mask method. Tabas et al. [Gauge] derive a differentiable bijection based on Minkowski functionals and apply it to power systems. Chen et al. [BoundaryProjectionLoss] define differentiable projection layers relying on convex constraints. While these approaches are, in principle, differentiable, previous work only utilises them to modify policy gradients.
2.3 IMPLICIT LAYERS
Defining the aforementioned safeguards can often not be done explicitly in closed form. Instead, they can only be formulated implicitly as a separate optimisation problem. Implicit layers [ImplicitLayers, CVXPY-Layers] enable an efficient backpropagation through the solution of this separate optimisation problem without unrolling the solver steps. They decouple the forward and backward pass and analytically differentiate via the implicit function theorem [ImplicitFunctionTheorem], which allows for constant training memory. Implicit layers are a potent paradigm that can be utilised for the tuning of controller parameters [CVXPY-Control], model identification [CVXPY-Models], and safeguarding [BoundaryProjectionLoss]. Given the potential complexity of general optimisation problems being NP-hard, it is crucial to approach the implicit formulation with diligence. If a restriction to convex cone programs is possible, solutions can be computed efficiently in polynomial time [diffcp], thereby facilitating a swift forward pass [GeometricProgramming][ConvexConicPrograms]. Moreover, this allows the utilisation of CVXPY [CVXPY, CVXPY2] to formulate the problem, which automatically picks an efficient solver and translates the problem to the desired solver formulation.
3 PRELIMINARIES
We use zonotopes to represent safe sets due to their compact representation and closedness under linear maps and Minkowski sums. They are convex, restricted polytopes and are defined as
$\mathcal{Z}=\left\{c+G\beta\,|\,\|{\beta}\|_{\infty}\leq 1\right\}=\left<c,G\right>\,$ (1)
with centre $c\in\mathbb{R}^{d}$, generator matrix $G\in\mathbb{R}^{d\times n}$, and scaling factors $\beta\in\mathbb{R}^{n}$. Zonotopes with orthogonal generators and $d=n$ are boxes which we use as representations for feasible and noise sets. We utilise the following properties of zonotopes to formulate our safeguards. The Minkowski sum of two zonotopes $\mathcal{Z}_{1},\mathcal{Z}_{2}\subset\mathbb{R}^{d}$ is [RCI-Sets, Eq. 7a]
$\mathcal{Z}_{1}\oplus\mathcal{Z}_{2}=\left<c_{1}+c_{2},\begin{bmatrix}G_{1}&G_{2}\end{bmatrix}\right>\,.$ (2)
Translating a zonotope is equivalent to translating the centre. Linearly mapping by $M\in\mathbb{R}^{n\times d}$ yields [RCI-Sets, Eq. 7b]
$M\mathcal{Z}=\left<Mc,MG\right>\,.$ (3)
A support function of a set describes the farthest extent of the set in a given direction. The support function of a zonotope is [SupportFunction, Lemma 1]
$\rho_{\mathcal{Z}}(v)=v^{T}c+\left\lVert G^{T}v\right\rVert_{1}\,.$ (4)
A point $p\in\mathbb{R}^{d}$ is contained in a zonotope if [zonotope_point_containment, Eq. 6]
$1\geq\min_{\gamma\in\mathbb{R}^{n}}\|\gamma\|_{\infty}\,\text{s.t.}\,p=c+G\gamma\,.$ (5)
Determining the containment of a zonotope in another zonotope is co-NP complete [zonotope_point_containment] but a sufficient condition for $\mathcal{Z}_{1}\subseteq\mathcal{Z}_{2}$ is [zonotope_containment, Eq. 15]
$1\geq$ $\!\min_{\gamma\in\mathbb{R}^{n_{2}},\Gamma\in\mathbb{R}^{n_{2}\times n_{1}}}$ $\left\lVert\begin{bmatrix}\Gamma&\gamma\end{bmatrix}\right\rVert_{\infty}$ (6a)
subject to $G_{1}=G_{2}\Gamma$ (6b)
$c_{2}-c_{1}=G_{2}\gamma\,.$ (6c)
Both containment problems are linear.
4 PROBLEM STATEMENT
The class of problems we tackle can be described by an augmented Markov decision process $(\mathcal{S},\mathcal{A},\mathcal{A}_{s},P_{f},r)$ of the following elements: the feasible state set $\mathcal{S}\subset\mathbb{R}^{d_{\mathcal{S}}}$, the feasible action set $\mathcal{A}\subset\mathbb{R}^{d}$, the safe action set $\mathcal{A}_{s}\subseteq\mathcal{A}$, the transition probability density function $P_{f}(s_{i+1}|s_{i},a_{i})$, and the reward function $r(s_{i},a_{i},s_{i+1})=r_{i}$. Axis-aligned boxes represent the feasible state and action set and the noise set. For the safeguarding, we employ an overapproximative transition function $f(s_{i},a_{i},w_{i})=s_{i+1}$ with bounded noise $w_{i}\sim\mathcal{W}=\left<c_{\mathcal{W}},G_{\mathcal{W}}\right>\subset\mathbb{R}^{d_{\mathcal{S}}}$, which can, for example, be obtained using reachset-conformant identification [Reachability, Overapproximation, Overapproximation1]. The transition and reward functions are continuously differentiable. The safe action set is either given explicitly as a zonotope, induced by a safe state zonotope $\mathcal{S}_{s}$, or the intersection of both constraints. The induced safe action set is
$\mathcal{A}_{s}(s_{i})=\left\{a_{i}\mid a_{i}\in\mathcal{A},\mathcal{S}_{i+1}(a_{i},s_{i})\subseteq\mathcal{S}_{s}\right\}$ (7)
with the next state set
$\mathcal{S}_{i+1}=f(s_{i},a_{i},\mathcal{W})=\{f(s_{i},a_{i},w_{i})\mid w_{i}\in\mathcal{W}\}\,.$ (8)
We seek a safeguarded, stochastic policy that maximises the expected, discounted, cumulative reward over a finite horizon $N$
$\pi^{*}(a|s)=\!\operatorname*{argmax}_{\pi(a|s)}\operatorname*{\mathbb{E}}_{\begin{subarray}{c}a\sim\pi(a|s)\\ s_{i+1}\sim P_{f}\end{subarray}}\sum_{i=0}^{N}\delta^{i}\,r(s_{i},{a_{s}}_{i},s_{i+1})$ (9)
with the safe action $a_{s}$, the continuously differentiable safeguarding $g_{\mathcal{A}_{s}}(a_{i},s_{i})={a_{s}}_{i}$, and discount factor $\delta\in(0,1)$.
5 METHOD
We show the policy, safeguard and environment relations in Figure 1. The safeguards are applied after the policy output. To train the policy, we calculate the gradient of the reward with respect to the policy output
$\frac{\partial r_{i}}{\partial a_{i}}=\left(\frac{\partial r_{i}}{\partial{a_{s}}_{i}}+\frac{\partial r_{i}}{\partial s_{i+1}}\frac{\partial s_{i+1}}{\partial{a_{s}}_{i}}\right)\frac{\partial{a_{s}}_{i}}{\partial a_{i}}\,.$ (10)
Since the policy output depends on the previous state, full backpropagation requires unrolling the trajectory to determine how all previous policy outputs affect the current reward.
Figure 1: Description of the forward (top) and backward (bottom) pass of the provably safe analytic policy optimisation.
The following subsections describe how we safeguard the analytic policy optimisation. First, we formulate the required and desired properties of a safeguard in the aforementioned differentiable setting. Then, we introduce the two safeguards used in this work: boundary projection (BP) [BoundaryProjection] and ray mask (RM) [RayMasking].
5.1 REQUIRED AND DESIRED PROPERTIES
The required properties of a safeguard for analytic gradient-based reinforcement learning are the containment of its codomain in the safe set and the existence of a Clark generalised derivative [Clarke] everywhere, which implies that the safeguarding is at least class $C^{0}$. Beyond the required, there are additional desired properties. First, the safeguarding should be class $C^{1}$ and provide full-rank Jacobians $\frac{\partial a_{s}}{\partial a}$ everywhere to facilitate learning and prevent information loss. Second, the number of interventions by the safeguarding should be minimal throughout training and inference. Minimal interventions also serve the last desired property of fast computation. In summary, the safeguard must
  1. P1
    map any action to a safe action,
  2. P2
    be subdifferentiable everywhere, and therefore class $C^{0}$
and should
  1. P3
    be class $C^{1}$ and provide full-rank Jacobians everywhere, i.e. be a local diffeomorphism,
  2. P4
    intervene rarely, and
  3. P5
    compute quickly.
The two safeguards we present offer different trade-offs between the desired properties.
5.2 BOUNDARY PROJECTION
The boundary projection safeguard [BoundaryProjection] maps actions to the safe actions with the smallest distance. It maps only the required unsafe actions to the nearest boundary point in the safe action set. We show an exemplary mapping with boundary projection from an unsafe action $a$ to a safe action $a_{s}$ in Figure 2.
Figure 2: Boundary projection maps unsafe actions to the boundary of the safe action set.
The safe action is the solution to
$\!\min_{a_{s}}$ $\left\lVert a-a_{s}\right\rVert_{2}^{2}$ (11a)
subject to $a_{s}\in\mathcal{A}_{s}\,.$ (11b)
If the safe action set is explicitly available, Equation 5 provides the containment constraint, which yields a quadratic problem.
Induced safe action sets could yield non-convex constraints, making the problem significantly more expensive. However, if the transition function is linear with respect to the action and the noise,
$\frac{\partial^{2}f}{\partial a^{2}}$ $=0$ (12)
$\frac{\partial^{2}f}{\partial w^{2}}$ $=0\,,$
the first-order Taylor expansion of the continuously differentiable transition function
$f(s_{i},a_{i},w_{i})\approx f\bigg{|}_{lin}$ $+\frac{\partial f}{\partial s}\bigg{|}_{lin}(s_{i}-s_{lin})$ (13)
$+\frac{\partial f}{\partial a}\bigg{|}_{lin}(a_{i}-a_{lin})$
$+\frac{\partial f}{\partial w}\bigg{|}_{lin}(w_{i}-w_{lin})$
around the current state is exact, if evaluated only at that state. This Taylor expansion can be utilised to compute the next set of states
$\mathcal{S}_{i+1}(s_{i},a_{i})=\left<f(s_{i},a_{i},c_{\mathcal{W}}),\frac{\partial f}{\partial w}\bigg{|}_{lin}G_{\mathcal{W}}\right>\,,$ (14)
which allows a replacement of Equation 11b with $\mathcal{S}_{i+1}(a_{i})\subseteq\mathcal{S}_{s}$. This containment constraint is either again Equation 5 for $\mathcal{W}=\emptyset$ or Equation 6 yielding again a quadratic problem.
In Euclidean space, unsafe actions are mapped orthogonally to the boundary $\partial\mathcal{A}_{s}$ of the safe action set. The distance between the initial and mapped action decreases smoothly with the distance from the unsafe action to the boundary until it is zero for safe actions. However, the mapping location can change abruptly between unsafe actions on different sides of the edges of the safe set. Moreover, any boundary point $b_{\mathcal{A}_{S}}\in\partial\mathcal{A}_{s}$ is the solution to Equation 11 for all actions
$a=b_{\mathcal{A}_{S}}+t\cdot v\quad\forall t,\,\forall v$ (15)
with the normal, unit vectors $v\in\{v\mid v\cdot(a_{s}-b_{\mathcal{A}_{S}})\leq 0,\forall a_{s}\in\mathcal{A}_{s}\}\setminus\{0\}$ and a positive scalar $t$. Intuitively, a boundary point is the nearest safe action for all actions along a normal vector. Therefore, the safeguard cannot propagate gradient components in the mapping direction, such that
$\left(\frac{\partial r}{\partial a_{s}}\frac{\partial a_{s}}{\partial a}\right)^{T}v=0$ (16)
which is especially problematic for gradients parallel to $v$, which may happen if the optimal action is safe. In such a case, boundary projection eliminates the gradient, keeping the optimisation stuck indefinitely. It also means the Jacobian has a rank of at most $d-1$.
Lemma 1.
Let $\mathcal{A}_{s}$ be explicitly available with a non-empty interior, such that strong duality holds, and the number of generators is at least the action dimension ($n\geq d$). Then, for safe actions, the Jacobian of the boundary projection safeguard is the identity and thus has full rank. For unsafe actions, the Jacobian is a projection matrix of rank at most $d-1$.
{IEEEproof}
Equation 11 becomes
$\min_{a_{s},\gamma}$ $\left\lVert a-a_{s}\right\rVert_{2}^{2}$ (17a)
subject to $a_{s}=c_{\mathcal{A}_{s}}+G_{\mathcal{A}_{s}}\gamma$ (17b)
$\left\lVert\gamma\right\rVert_{\infty}\leq 1$ (17c)
with safe action set $\mathcal{A}_{s}=\left<c_{\mathcal{A}_{s}},G_{\mathcal{A}_{s}}\right>$. Translating Equation 17 to canonical, quadratic form yields the optimisation variable $z=\begin{bmatrix}a_{s}\\ \gamma\end{bmatrix}\in\mathbb{R}^{d+n}$ and the problem
$\min_{z}$ $\frac{1}{2}z^{T}Qz+q^{T}z$ (18a)
subject to $Az=b$ (18b)
$Kz\leq h$ (18c)
with
$Q$ $=\begin{bmatrix}2I_{d}&\textbf{0}\\ \textbf{0}&\textbf{0}\\ \end{bmatrix}\in\mathbb{R}^{(d+n)\times(d+n)}$ (19)
$q$ $=-2\begin{bmatrix}a\\ \textbf{0}\end{bmatrix}\in\mathbb{R}^{d+n}$ (20)
$A$ $=\begin{bmatrix}I_{d}&-G_{\mathcal{A}_{s}}\end{bmatrix}\in\mathbb{R}^{d\times(d+n)}$ (21)
$b$ $=c_{\mathcal{A}_{s}}\in\mathbb{R}^{d}$ (22)
$K$ $=\begin{bmatrix}\textbf{0}&I_{n}\end{bmatrix}\in\mathbb{R}^{n\times(d+n)}$ (23)
$h$ $=\textbf{1}\in\mathbb{R}^{n}\,,$ (24)
where the subscript of the identity denotes its size, and bold scalars are matrices of appropriate size filled with the scalar. To compute the rank of the Jacobian, we start from the differentials of the KKT conditions [DifferentiateQuadraticProgram, Eq. 6]
$\begin{bmatrix}Q&K^{T}&A^{T}\\ \lambda_{D}^{*}K&Kz_{D}^{*}-h_{D}&\textbf{0}\\ A&\textbf{0}&\textbf{0}\end{bmatrix}\begin{bmatrix}\mathop{}\!\mathrm{d}z\\ \mathop{}\!\mathrm{d}\lambda\\ \mathop{}\!\mathrm{d}\nu\end{bmatrix}=$ (25)
$\begin{bmatrix}-\mathop{}\!\mathrm{d}Qz^{*}-\mathop{}\!\mathrm{d}q-\mathop{}\!\mathrm{d}K^{T}\lambda^{*}-\mathop{}\!\mathrm{d}A^{T}\nu^{*}\\ -\lambda_{D}^{*}\mathop{}\!\mathrm{d}Kz^{*}+\lambda_{D}^{*}\mathop{}\!\mathrm{d}h\\ -\mathop{}\!\mathrm{d}Az^{*}+\mathop{}\!\mathrm{d}b\end{bmatrix}$
where the subscript $D$ denotes a diagonalised vector, the superscript * optimal values, $\nu\in\mathbb{R}^{d}$ the dual variables on the equality constraints, $\lambda\in\mathbb{R}^{n}$ the dual variables on the inequality constraints, and $\mathop{}\!\mathrm{d}$ a differential. To obtain the Jacobian with respect to the action, we substitute $\mathop{}\!\mathrm{d}q=\frac{\mathop{}\!\mathrm{d}q}{\mathop{}\!\mathrm{d}a}\mathop{}\!\mathrm{d}a=-2\begin{bmatrix}I_{d}\\ \textbf{0}\end{bmatrix}\mathop{}\!\mathrm{d}a$ and all other differential terms with zero leaving
$\begin{bmatrix}Q&K^{T}&A^{T}\\ \lambda_{D}^{*}K&Kz_{D}^{*}-h_{D}&\textbf{0}\\ A&\textbf{0}&\textbf{0}\end{bmatrix}\begin{bmatrix}\frac{\mathop{}\!\mathrm{d}z}{\mathop{}\!\mathrm{d}a}\\ \frac{\mathop{}\!\mathrm{d}\lambda}{\mathop{}\!\mathrm{d}a}\\ \frac{\mathop{}\!\mathrm{d}\nu}{\mathop{}\!\mathrm{d}a}\end{bmatrix}=\begin{bmatrix}2\begin{bmatrix}I_{d}\\ \textbf{0}\end{bmatrix}\\ \textbf{0}\\ \textbf{0}\end{bmatrix}\,.$ (26)
We insert the variables and describe the system in expanded form as
$\begin{bmatrix}2I_{d}&\textbf{0}&\textbf{0}&I_{d}\\ \textbf{0}&\textbf{0}&I_{n}&-G_{\mathcal{A}_{s}}^{T}\\ \textbf{0}&\lambda_{D}^{*}&\gamma_{D}^{*}-I_{n}&\textbf{0}\\ I_{d}&-G_{\mathcal{A}_{s}}&\textbf{0}&\textbf{0}\end{bmatrix}\begin{bmatrix}\frac{\mathop{}\!\mathrm{d}a_{s}}{\mathop{}\!\mathrm{d}a}\\ \frac{\mathop{}\!\mathrm{d}\gamma}{\mathop{}\!\mathrm{d}a}\\ \frac{\mathop{}\!\mathrm{d}\lambda}{\mathop{}\!\mathrm{d}a}\\ \frac{\mathop{}\!\mathrm{d}\nu}{\mathop{}\!\mathrm{d}a}\\ \end{bmatrix}=\begin{bmatrix}2I_{d}\\ \textbf{0}\\ \textbf{0}\\ \textbf{0}\\ \end{bmatrix}\,.$ (27)
This yields a coupled system of matrix equations
$2\frac{\mathop{}\!\mathrm{d}a_{s}}{\mathop{}\!\mathrm{d}a}+\frac{\mathop{}\!\mathrm{d}\nu}{\mathop{}\!\mathrm{d}a}$ $=2I_{d}$ (28)
$\frac{\mathop{}\!\mathrm{d}\lambda}{\mathop{}\!\mathrm{d}a}-G_{\mathcal{A}_{s}}^{T}\frac{\mathop{}\!\mathrm{d}\nu}{\mathop{}\!\mathrm{d}a}$ $=\textbf{0}$ (29)
$\lambda_{D}^{*}\frac{\mathop{}\!\mathrm{d}\gamma}{\mathop{}\!\mathrm{d}a}+(\gamma_{D}^{*}-I_{n})\frac{\mathop{}\!\mathrm{d}\lambda}{\mathop{}\!\mathrm{d}a}$ $=\textbf{0}$ (30)
$\frac{\mathop{}\!\mathrm{d}a_{s}}{\mathop{}\!\mathrm{d}a}-G_{\mathcal{A}_{s}}\frac{\mathop{}\!\mathrm{d}\gamma}{\mathop{}\!\mathrm{d}a}$ $=\textbf{0}\,.$ (31)
We solve this system by substitution starting from Equation 31
$\frac{\mathop{}\!\mathrm{d}a_{s}}{\mathop{}\!\mathrm{d}a}=G_{\mathcal{A}_{s}}\frac{\mathop{}\!\mathrm{d}\gamma}{\mathop{}\!\mathrm{d}a}\,,$ (32)
which we substitute into Equation 28
$\frac{\mathop{}\!\mathrm{d}\nu}{\mathop{}\!\mathrm{d}a}=2I_{d}-2G_{\mathcal{A}_{s}}\frac{\mathop{}\!\mathrm{d}\gamma}{\mathop{}\!\mathrm{d}a}\,.$ (33)
We substitute this into Equation 29
$\frac{\mathop{}\!\mathrm{d}\lambda}{\mathop{}\!\mathrm{d}a}=2G_{\mathcal{A}_{s}}^{T}-2G_{\mathcal{A}_{s}}^{T}G_{\mathcal{A}_{s}}\frac{\mathop{}\!\mathrm{d}\gamma}{\mathop{}\!\mathrm{d}a}\,,$ (34)
with which we decouple Equation 30
$(\lambda_{D}^{*}-2(\gamma_{D}^{*}-I_{n})G_{\mathcal{A}_{s}}^{T}G_{\mathcal{A}_{s}})\frac{\mathop{}\!\mathrm{d}\gamma}{\mathop{}\!\mathrm{d}a}=-2(\gamma_{D}^{*}-I_{n})G_{\mathcal{A}_{s}}^{T}\,.$ (35)
To solve this linear system, we utilise the KKT complementarity slackness conditions in element-wise notation
$\lambda_{i}^{*}(K_{i,:}z_{*}-h_{i})=\lambda_{i}^{*}(\gamma_{i}^{*}-1)=0\quad\forall i=1,\ldots,n\,.$ (36)
The constraint is inactive $\lambda_{i}^{*}=0$ or active $\gamma_{i}^{*}=1$ to fulfil the conditions. We define the index set of active constraints as $\mathcal{I}_{a}=\left\{i\mid\gamma_{i}^{*}=1\right\}$ and of inactive constraints as $\mathcal{I}_{i}=\left\{i\mid\lambda_{i}^{*}=0\right\}$. We examine the rows of Equation 35
$\lambda_{i}^{*}\left(\frac{\mathop{}\!\mathrm{d}\gamma}{\mathop{}\!\mathrm{d}a}\right)_{i,:}-2(\gamma_{i}^{*}-1)(G_{\mathcal{A}_{s}}^{T}G_{\mathcal{A}_{s}})_{i,:}\frac{\mathop{}\!\mathrm{d}\gamma}{\mathop{}\!\mathrm{d}a}$ (37)
$=-2(\gamma_{i}^{*}-1)(G_{\mathcal{A}_{s}}^{T})_{i,:}\,,$
where we utilise the diagonalised form of the optimal variables and denote full rows or columns with a colon subscript. Next, we combine this with the index sets
$\lambda_{i}^{*}\left(\frac{\mathop{}\!\mathrm{d}\gamma}{\mathop{}\!\mathrm{d}a}\right)_{i,:}$ $=0$ $i\in\mathcal{I}_{a}$ (38)
$(G_{\mathcal{A}_{s}}^{T}G_{\mathcal{A}_{s}})_{i,:}\frac{\mathop{}\!\mathrm{d}\gamma}{\mathop{}\!\mathrm{d}a}$ $=(G_{\mathcal{A}_{s}}^{T})_{i,:}$ $i\in\mathcal{I}_{i}\,.$ (39)
We partition the Jacobian according to the index sets $\frac{\mathop{}\!\mathrm{d}\gamma}{\mathop{}\!\mathrm{d}a}=\begin{bmatrix}\left(\frac{\mathop{}\!\mathrm{d}\gamma}{\mathop{}\!\mathrm{d}a}\right)_{\mathcal{I}_{a},:}\\ \left(\frac{\mathop{}\!\mathrm{d}\gamma}{\mathop{}\!\mathrm{d}a}\right)_{\mathcal{I}_{i},:}\end{bmatrix}$. The assumption of strong duality implies strict complementarity $\lambda_{i}^{*}>0$. Therefore, Equation 38 specifies the submatrix $\left(\frac{\mathop{}\!\mathrm{d}\gamma}{\mathop{}\!\mathrm{d}a}\right)_{\mathcal{I}_{a},:}=\textbf{0}$. The left-hand side of Equation 39 reduces to
$(G_{\mathcal{A}_{s}}^{T}G_{\mathcal{A}_{s}})_{i,:}\frac{\mathop{}\!\mathrm{d}\gamma}{\mathop{}\!\mathrm{d}a}$ $=\sum_{j}{(G_{\mathcal{A}_{s}}^{T}G_{\mathcal{A}_{s}})_{i,j}\left(\frac{\mathop{}\!\mathrm{d}\gamma}{\mathop{}\!\mathrm{d}a}\right)_{j,:}}$ (40)
$=\sum_{j\in\mathcal{I}_{i}}{(G_{\mathcal{A}_{s}}^{T}G_{\mathcal{A}_{s}})_{i,j}\left(\frac{\mathop{}\!\mathrm{d}\gamma}{\mathop{}\!\mathrm{d}a}\right)_{j,:}}$
$=\sum_{j\in\mathcal{I}_{i}}{\left(\left(G_{\mathcal{A}_{s}}^{T}\right)_{i,:}\left(G_{\mathcal{A}_{s}}\right)_{:,j}\right)\left(\frac{\mathop{}\!\mathrm{d}\gamma}{\mathop{}\!\mathrm{d}a}\right)_{j,:}}\,,$
using Equation 38. We isolate the remaining submatrix
$\left(\left(G_{\mathcal{A}_{s}}^{T}\right)_{\mathcal{I}_{i},:}\left(G_{\mathcal{A}_{s}}\right)_{:,\mathcal{I}_{i}}\right)\left(\frac{\mathop{}\!\mathrm{d}\gamma}{\mathop{}\!\mathrm{d}a}\right)_{\mathcal{I}_{i},:}=(G_{\mathcal{A}_{s}}^{T})_{\mathcal{I}_{i},:}\,.$ (41)
The system is always solvable since $n\geq d$, and the right-hand side lies trivially in the column space of the left-hand side. We utilise the Moore-Penrose inverse to solve the system
$\left(\frac{\mathop{}\!\mathrm{d}\gamma}{\mathop{}\!\mathrm{d}a}\right)_{\mathcal{I}_{a},:}$ $=\textbf{0}$ (42)
$\left(\frac{\mathop{}\!\mathrm{d}\gamma}{\mathop{}\!\mathrm{d}a}\right)_{\mathcal{I}_{i},:}$ $=\left(\left(G_{\mathcal{A}_{s}}^{T}\right)_{\mathcal{I}_{i},:}\left(G_{\mathcal{A}_{s}}\right)_{:,\mathcal{I}_{i}}\right)^{\dagger}\left(G_{\mathcal{A}_{s}}^{T}\right)_{\mathcal{I}_{i},:}\,.$ (43)
We plug this result into Equation 32
$\frac{da_{s}}{da}$ $=\begin{bmatrix}\left(G_{\mathcal{A}_{s}}\right)_{:,\mathcal{I}_{a}}&\left(G_{\mathcal{A}_{s}}\right)_{:,\mathcal{I}_{i}}\end{bmatrix}\begin{bmatrix}\left(\frac{\mathop{}\!\mathrm{d}\gamma}{\mathop{}\!\mathrm{d}a}\right)_{\mathcal{I}_{a},:}\\ \left(\frac{\mathop{}\!\mathrm{d}\gamma}{\mathop{}\!\mathrm{d}a}\right)_{\mathcal{I}_{i},:}\end{bmatrix}$ (44)
$=\left(G_{\mathcal{A}_{s}}\right)_{:,\mathcal{I}_{i}}\left(\frac{d\gamma}{da}\right)_{\mathcal{I}_{i},:}$
$=\left(G_{\mathcal{A}_{s}}\right)_{:,\mathcal{I}_{i}}\left(\left(G_{\mathcal{A}_{s}}^{T}\right)_{\mathcal{I}_{i},:}\left(G_{\mathcal{A}_{s}}\right)_{:,\mathcal{I}_{i}}\right)^{\dagger}\left(G_{\mathcal{A}_{s}}^{T}\right)_{\mathcal{I}_{i},:}$
The resulting matrix is the projection matrix onto the column space of $\left(G_{\mathcal{A}_{s}}\right)_{:,\mathcal{I}_{i}}$. It is also unique since any homogeneous part of the solution lies in the null space of $\left(G_{\mathcal{A}_{s}}\right)_{:,\mathcal{I}_{i}}$, see Section 7.6 for details.
The rank of a projection matrix is equal to the design matrix itself $\text{rank}(\frac{\mathop{}\!\mathrm{d}a_{s}}{\mathop{}\!\mathrm{d}a})=\text{rank}(\left(G_{\mathcal{A}_{s}}\right)_{:,\mathcal{I}_{i}})$. Due to the size of $G_{\mathcal{A}_{s}}$, the rank of the Jacobian is at most $d$ when all inequality constraints are inactive, which is only the case for interior points of the zonotope, i.e. safe actions. However, for boundary points, i.e. mapped unsafe actions, at least one constraint is active. Therefore, the rank is at most $d-1$. To regain a gradient component in the mapping direction, we augment the policy loss function $l_{r}(a_{s},s)$ with a regularisation term [BoundaryProjectionLoss, Eq. 16]
$l(a,s,a_{s})=l_{r}(a_{s},s)+c_{d}\left\lVert a_{s}-a\right\rVert_{2}^{2}\,,$ (45)
which augments the gradient by $\frac{\partial l}{\partial a}=\frac{\partial l_{r}}{\partial a}+2c_{d}(a_{s}-a)(\frac{\partial a_{s}}{\partial a}-I)$. In addition to gradient augmentation, regularisation encourages the policy to favour safe actions from the start, which is desirable for Property P4. The coefficient $c_{d}$ scales the regularisation to remain small relative to the reward-informed loss $l_{r}(a_{s},s)$ yet large enough to produce a meaningful gradient component in the mapping direction. If $c_{d}$ is too large, the gradient may become uninformative or vanish for unsafe actions. Vanishing may occur with action spaces normalised with a hyperbolic tangent function, as it saturates for exceedingly large values.
5.3 RAY MASK
The ray mask technique [RayMasking] maps every action radially towards the centre of the safe action set, as shown in Figure 3.
Figure 3: The Ray mask maps actions towards the safe centre in proportion to the relative safety domain length.
The safe action is [RayMasking, Eq. 6]
$a_{s}=c_{\mathcal{A}_{s}}+\frac{\lambda_{\mathcal{A}_{s}}}{\lambda_{\mathcal{A}}}(a-c_{\mathcal{A}_{s}})$ (46)
with the centre of the safe action set $c_{\mathcal{A}_{s}}$, $\lambda_{\mathcal{A}_{s}}=\left\lVert c_{\mathcal{A}_{s}}-b_{\mathcal{A}_{s}}\right\rVert_{2}$, and $\lambda_{\mathcal{A}}=\left\lVert c_{\mathcal{A}_{s}}-b_{\mathcal{A}}\right\rVert_{2}$, where $b_{\mathcal{A}_{s}}\in\partial\mathcal{A}_{s}$ is the safe boundary point and $b_{\mathcal{A}}\in\partial A$ the feasible boundary point along the ray. The feasible boundary can be explicitly computed when the feasible action set is an axis-aligned box.
For explicitly available safe action zonotopes, the safe centre is defined as the centre of the zonotope, and the safe boundary is the solution to
$\!\max_{\alpha}$ $\alpha$ (47a)
subject to $b_{\mathcal{A}_{s}}=c_{\mathcal{A}_{s}}+\alpha(a-c_{\mathcal{A}_{s}})$ (47b)
$b_{\mathcal{A}_{s}}\in\mathcal{A}_{s}$ (47c)
with the same containment constraint Equation 5 yielding a linear problem.
For induced safe action sets, the same as for boundary projection applies: the problem becomes more expensive to solve; however, under Equation 12, the containment constraint can be replaced, which yields a linear problem again. Additionally, ray masking requires a star-shaped safe action set to ensure that the centre and line segment from the boundary point to the centre lie within the set. All convex sets are star-shaped. While the induced safe action set is not necessarily convex for arbitrary nonlinear transition functions, it is convex under Equation 12.
Lemma 2.
Let $\mathcal{A}_{s}$ be induced by a zonotope with a transition function that is linear with respect to the action and the noise, i.e. Equation 12 holds. Then, the induced safe action set is convex.
{IEEEproof}
As the noise set $\mathcal{W}$ is a box, we use Equation 3 and Equation 13 to derive the next set of states
$\mathcal{S}_{i+1}(a_{i})=\left<\text{const.},\frac{\partial f}{\partial w}\bigg{|}_{lin}G_{\mathcal{W}}\right>+\frac{\partial f}{\partial a}\bigg{|}_{lin}a_{i}\,.$ (48)
with all terms independent of $a_{i}$ absorbed into const. to shorten notation. Placing this into Equation 7 and abbreviating further yields
$\mathcal{A}_{s}(s_{i})=\left\{a_{i}\in\mathcal{A}:\mathcal{Z}_{const.}+\frac{\partial f}{\partial a}\bigg{|}_{lin}a_{i}\subseteq\mathcal{S}_{s}\right\}\,.$ (49)
Since applying the matrix $\frac{\partial f}{\partial a}\bigg{|}_{lin}$ is a linear operation and convexity preserving, the safe action set is convex if and only if the set of all translations $\mathcal{T}=\left\{t:\mathcal{Z}_{const.}+t\subseteq\mathcal{S}_{s}\right\}$ is convex. The definition of convexity dictates
$\lambda x+(1-\lambda)y\in\mathcal{S}_{s}$ (50)
for any $x,y\in\mathcal{S}_{s}$ and $\lambda\in\left[0,1\right]$. It also applies to $\mathcal{T}$, since for any $t_{1},t_{2}\in\mathcal{T}$ setting $x=\mathcal{Z}_{const.}+t_{1}$ and $y=\mathcal{Z}_{const.}+t_{2}$ yields
$\lambda(\mathcal{Z}_{const.}+t_{1})+(1-\lambda)(\mathcal{Z}_{const.}+t_{2})$ (51)
$=\mathcal{Z}_{const.}+(\lambda t_{1}+(1-\lambda)t_{2})\in\mathcal{S}_{s}\,,$
which is the respective convexity condition of $\mathcal{T}$. Moreover, ray masking also requires $c_{\mathcal{A}_{s}}$, which is not explicitly available in induced safe action sets. We present two approaches to approximate it: orthogonal and zonotopic approximation. They are visualised in Figure 4.
Figure 4: Approximation of the safe centre by expanding a zonotope (left) and by orthogonally projecting through the boundary (right).
The zonotopic approach directly approximates the safe action set. The underapproximated zonotope $\mathcal{Z}_{\mathcal{A}_{s}}$ is the solution to
$\!\max_{c_{\mathcal{A}_{s}},l_{s}}$ $n]{\prod_{i=1}^{n}l_{s,i}}$ (52a)
subject to $\mathcal{Z}_{\mathcal{A}_{s}}=\left<c_{\mathcal{A}_{s}},G_{\mathcal{A}_{s}}\text{diag}(l_{s})\right>$ (52b)
$\mathcal{Z}_{\mathcal{A}_{s}}\subseteq\mathcal{A}$ (52c)
$\mathcal{S}_{i+1}=f(s_{i},\mathcal{Z}_{\mathcal{A}_{s}},\mathcal{W})$ (52d)
$\mathcal{S}_{i+1}\subseteq\mathcal{S}_{s}$ (52e)
with $n$ uniformally sampled generator directions $G_{\mathcal{A}_{s}}$. Since we seek a tight approximation, the objective is maximising the geometric mean of the generator lengths, which we use as a computationally cheaper proxy for the volume [RayMasking]. Generally, Equation 52d may be a non-convex constraint, yielding an expensive problem. However, under Equation 12 and using Equation 2 $\mathcal{S}_{i+1}$ is the zonotope
$\mathcal{S}_{i+1}=\left<f(s_{i},c_{\mathcal{W}}),\begin{bmatrix}\frac{\partial f}{\partial a}\bigg{|}_{lin}G_{\mathcal{Z}_{\mathcal{A}_{s}}}&\frac{\partial f}{\partial w}\bigg{|}_{lin}G_{\mathcal{W}}\end{bmatrix}\right>\,.$ (53)
This problem is conic, which makes it comparatively expensive to compute. The computation time also increases with the number of generators. Generally, the number of generators should be in the order of the number of dimensions to provide a good approximation but not too large since the geometric mean favours similar lengths.
The orthogonal approximation does not directly approximate the safe action set. Instead, it first approximates the safe boundary point $b_{\mathcal{A}_{s}}$ by the boundary projection technique. Then, it extends the orthogonal direction according to Equation 47 until hitting a second safe boundary point $b_{\mathcal{A}_{s,2}}$. The middle point between the two boundary points is the safe centre
$c_{\mathcal{A}_{s}}=\frac{b_{\mathcal{A}_{s}}+b_{\mathcal{A}_{s,2}}}{2}\,.$ (54)
Since the boundary projection technique only maps unsafe actions, the orthogonal approximation can only be used for unsafe actions. Compared to the zonotopic approximation, the orthogonal approximation requires the solution of one quadratic and one linear problem if the action is unsafe; otherwise, it requires only the quadratic problem to determine the safety.
The ray mask approach, excluding approximations, is smooth almost everywhere except for the safe set edges. In contrast to boundary projection, its Jacobian is of full rank.
Lemma 3.
Let $\mathcal{A}_{s}$ be convex. Then, the Jacobian of the ray mask has full rank.
{IEEEproof}
The easiest application and differentiation of the ray mask is in spherical coordinates, centred at the safe centre. We transform the coordinates with
$a_{o}=\begin{bmatrix}a_{o,r}\\ a_{o,1}\\ \vdots\\ a_{o,n-1}\end{bmatrix}=\text{spherical}(a-c_{\mathcal{A}_{s}})\,,$ (55)
where we adopt the convention of the radius being the first coordinate. In these coordinates, the ray mask modifies a single coordinate
$a_{s,o}=\begin{bmatrix}\frac{\lambda_{\mathcal{A}_{s}}}{\lambda_{\mathcal{A}}}a_{o,r}\\ a_{o,1}\\ \vdots\\ a_{o,n-1}\end{bmatrix}\,.$ (56)
Finally, we transform the safe action back
$a_{s}=\text{spherical}(a_{s,o})^{-1}+c_{\mathcal{A}_{s}}\,.$ (57)
The chain rule provides the Jacobian of the ray mask
$\frac{\partial a_{s}}{\partial a}=\frac{\partial a_{s}}{\partial a_{s,o}}\frac{\partial a_{s,o}}{\partial a_{o}}\frac{\partial a_{o}}{\partial a}\,,$ (58)
where the inverse function theorem [ImplicitFunctionTheorem] relates the Jacobians of the transformations as
$\frac{\partial a_{s}}{\partial a_{s,o}}=\frac{\partial a_{o}}{\partial a}^{-1}\,.$ (59)
This relation characterises a similarity transformation, which means $\frac{\partial a_{s}}{\partial a}$ and $\frac{\partial a_{s,o}}{\partial a_{o}}$ are similar and share the same eigenvalues. The Jacobian $\frac{\partial a_{s,o}}{\partial a_{o}}$ is
[…]
which is triangular and, therefore, has the eigenvalues in the diagonal elements
$\sigma_{i}=\begin{cases}\frac{\lambda_{\mathcal{A}_{s}}}{\lambda_{\mathcal{A}}}&\text{if }i=1\\ 1&\text{else}\,.\end{cases}$ (61)
Since they are all non-zero and the Jacobian is a square matrix, it has full rank. While the ray mask propagates gradient components in the mapping direction, they are still diminished. This reduction is most obvious in the scenario, where the feasible and safe action set are spheres with coinciding centres and radii $r_{\mathcal{A}}>r_{\mathcal{A}_{s}}$, and the coordinate system is already spherical and centred. In this scenario, the Jacobian in Equation 60 reduces to
$\frac{\partial a_{s}}{\partial a}=\begin{bmatrix}\frac{r_{\mathcal{A}}}{r_{\mathcal{A}_{s}}}&\textbf{0}\\ \textbf{0}&I\end{bmatrix}\,$ (62)
which has a trivial eigenspace, as the Jacobian is diagonal. Consequently, the upstream gradient is only modified in the mapping direction by the factor $\frac{r_{\mathcal{A}}}{r_{\mathcal{A}_{s}}}<1$. We summarise the properties of the safeguards in Table 1.
We propose three possible modifications to the ray mask to improve its learning properties. First, we can increase the gradient in the mapping direction with the same regularisation term as in boundary projection, see Equation 45. Second, we can replace the Jacobian with an identity matrix. The gradient directions would still be correct if the optimal action is safe. However, the safety of the optimal action is generally unknown and depends on the environment. For unsafe optimal actions, the point of convergence would be the safe boundary point on the line $\overline{c_{\mathcal{A}_{s}}a_{opt}}$, which is no longer optimal. In the unaltered ray mask, the mapping distance decreases linearly with the distance to the safe centre, using Equation 46
$\left\lVert a_{s}-a\right\rVert_{2}=\left\lVert c_{\mathcal{A}_{s}}-a\right\rVert_{2}(1-\frac{\lambda_{\mathcal{A}_{s}}}{\lambda_{\mathcal{A}}})\,,$ (63)
where we use the fact that the problem always reduces to one dimension along the mapping direction. This slow decay often indicates interventions, as all actions far from the centre are substantially altered. Therefore, we propose a hyperbolic mapping
$a_{s}=c_{\mathcal{A}_{s}}+(a-c_{\mathcal{A}_{s}})\frac{\lambda_{\mathcal{A}_{s}}}{\lambda_{a}}\frac{\tanh{\frac{\lambda_{a}}{\lambda_{\mathcal{A}_{s}}}}}{\tanh{\frac{\lambda_{\mathcal{A}}}{\lambda_{\mathcal{A}_{s}}}}}$ (64)
with $\lambda_{a}=\left\lVert a-c_{\mathcal{A}_{s}}\right\rVert$, which maps unsafe actions close to the boundary and safe actions close to themselves. The mapping distance decreases approximately linearly for unsafe actions but exponentially for safe actions. As it behaves similarly to boundary projection for unsafe actions, it suffers from the same gradient loss and benefits from a regularisation term.
Table 1: Properties of the unaltered safeguards.
Boundary Projection Ray Mask
Property P1 Safety
Property P2 Subdifferentiable
Property P3 Class $C^{1}$ $C^{1}$ almost everywhere
Full-rank Jacobians $d-1$
Property P4 Rare interventions only unsafe actions all actions
Property P5 Fast computation
For $\mathcal{A}_{s,explicit}$ 1 Quadratic Problem 1 Linear Problem
For $\mathcal{S}_{s}$ Convex 1 Quadratic Problem 0-1 Conic & 0-1 Quadratic & 1 Linear Problem
6 NUMERICAL EXPERIMENTS
This section tests our two main hypotheses:
  1. H1
    When considering safety guarantees, analytic gradient-based reinforcement learning outperforms sampling-based reinforcement learning.
  2. H2
    While existing differentiable safeguards can impede gradient flow and learning performance in analytic policy optimisation, our modifications successfully mitigate these issues, achieving equal or exceeding performance compared to unsafe training.
In the following subsections, we introduce our experimental setup, discuss the main hypotheses and provide additional insights.
6.1 SETUP
We conducted all experiments using ten different random seeds. Hyperparameters were tuned exclusively for non-safeguarded training and remained consistent throughout the safeguarded experiments [Optuna]. We assessed the quality of the final policy by calculating the mean episodic reward (Reward) achieved over a representative evaluation set alongside a 95% confidence interval computed using bootstrapping. For our convergence metrics, we tracked the number of steps until convergence (# Steps), defined as reaching within 5% of the final policy, and the number of runs that did not converge within the maximum allowed simulation steps (# Stuck). We excluded non-convergent runs from the reward and step calculations to ensure clarity.
We vary all three stages of the learning pipeline: policy optimisation algorithm, safeguarding, and simulation.
6.1.1 POLICY OPTIMISATION ALGORITHMS
We choose first-order SHAC [SHAC] as a learning algorithm over its successor Adaptive-Horizon Actor-Critic [AHAC] due to its maturity, stable convergence, and the lack of stiff dynamics in our tasks. We compare it with two sampling-based reinforcement learning algorithms: on-policy Proximal Policy Optimisation (PPO) [PPO] and off-policy Soft Actor-Critic (SAC) [SAC].
Because sampling-based algorithms do not backpropagate through the reward, they are not automatically informed about the safeguarding. Moreover, replacing unsafe actions with safe actions also poses problems for stochastic policies, which rely on the probabilities of the actions. In this work, we implement safeguarding as a post-processing step to the policy output without explicitly informing the sampling-based learning processes. This requires them to learn the dynamics associated with the safeguarding measures.
6.1.2 SAFEGUARDS
We evaluate the base versions of the boundary projection and ray mask as safeguards, where the zonotopic approach approximates the safe centre. We also evaluate all modifications to the safeguards individually and the combination of regularisation and the hyperbolic ray mask. We did not investigate all combinations of modifications due to computational constraints.
6.1.3 SIMULATIONS
We build our differentiable simulations according to the gymnasium framework [Gymnasium] and differentiate through the dynamics using PyTorch’s auto-differentiation engine [Pytorch]. We formulate the convex optimisation problems with CVXPY [CVXPY, CVXPY2] and backpropagate through them with CVXPYLayers [CVXPY-Layers]. Since we use vectorised environments to speed up training, early truncation is only possible for the entire vector. We select two differentiable simulations of classical control problems for our evaluation: balancing a pendulum and balancing a quadrotor [Quadrotor]. Their detailed descriptions are in Section 7.1.
6.2 EVALUATION OF POLICY OPTIMISATION ALGORITHMS
In this subsection, we evaluate Hypothesis H1, which was based on the superior performance in unsafe training and the natural integration of safeguards.
To establish the former, we first compare the learning algorithms in unsafe training. We display the key metrics in Table 2 and the learning curves in Section 7.2. SHAC converged to the best policies across both environments, whereas SAC performed slightly and PPO substantially worse. Notable exceptions are the slightly faster convergence speed of SAC in the pendulum environment and non-convergent runs for SHAC in the pendulum environment and SAC in the quadrotor environment.
Table 2: Comparison of learning algorithms in unsafe training.
Pendulum
Algorithm # Steps Reward # Stuck
SHAC 15360 -8.780 1
PPO 81600 -596.4 0
SAC 12520 -14.07 0
Quadrotor
Algorithm # Steps Reward # Stuck
SHAC 38528 -157.5 0
PPO 134400 -1710 0
SAC 105168 -242.2 2
After establishing the superior performance in unsafe training, we proceed to the hypothesis, comparing safeguarded training. We show the key metrics in Table 3 and the learning curves in Section 7.3 and Section 7.4. Our observations support the hypothesis since the final policy and convergence speed of SHAC were superior in all experiments except for the early pendulum policies of SAC. There SAC initially reached near-optimal performance within the first evaluation but later diverged. Under uninformed safeguarding, SAC should benefit from its off-policy nature, but the reliance on the probability of the chosen action outweighs this effect. This issue is most noticeable in the critic loss, which is continuously divergent. The reported lower number of steps for SAC in the pendulum environment is an artefact of the inferior policy. On the quadrotor task, Soft Actor-Critic learns ineffectively until the buffer resets roughly twice, at which point a jump in performance is visible consistently. The poor initial performance could result from uninformative earlier samples, although the underlying reason for the drastic performance increase is unclear. PPO benefitted from the guided exploration in the selected environments, as safety is strongly tied to reward, but did not achieve the same sample efficiency as SHAC. The non-convergent runs of SHAC with boundary projection can be attributed to a total loss of gradient information as outlined in Equation 16 since the action space in the pendulum environment is one-dimensional.
Table 3: Comparison of learning algorithms in safeguarded training.
Pendulum
Safe- guard Algorithm # Steps Reward # Stuck
BP SHAC 25600 -8.402 2
PPO 81600 -10.34 0
SAC 2504 -1083 0
RM SHAC 46080 -8.412 0
PPO 81600 -12.19 0
SAC 2504 -424.1 0
Quadrotor
Safe- guard Algorithm # Steps Reward # Stuck
BP SHAC 82560 -333.6 0
PPO 156800 -402.9 0
SAC 100160 -338.0 0
RM SHAC 110080 -251.7 0
PPO 156800 -379.3 0
SAC 95152 -377.9 0
6.3 EVALUATION OF SAFEGUARDS
Next, we evaluate Hypothesis H2 by comparing the various safeguards to unsafe training on SHAC, with the aggregated learning curves visualised in Figure 5 and the number of non-convergent runs in Table 4. Compared to unsafe training, we expect performance to decline for the unaltered (base) safeguards when the optimal action is safe, as is the case for most actions in balancing scenarios. The impact should be more severe for the unaltered boundary projection than the ray mask, attributed to the lack of gradient propagation in the mapping direction.
Figure 5: Comparison of SHAC in unsafe and safeguarded training via boundary projection (top) and ray mask (bottom).
Table 4: Number of non-convergent runs for the various safeguards.
Pendulum
Modification # Stuck
BP 2
BP Regularisation 1
RM 0
RM Regularisation 0
RM Passthrough Gradient 2
RM Hyperbolic 1
RM Hyperbolic and Regularisation 0
Quadrotor
Modification # Stuck
BP 0
BP Regularisation 0
RM 0
RM Regularisation 0
RM Passthrough Gradient 1
RM Hyperbolic 0
RM Hyperbolic and Regularisation 0
The impeded learning performance was visible in the quadrotor environment as a drop in policy quality and convergence speed for boundary projection, as the agents were learning very slowly or stalled for several individual runs completely. We made the same observation for the ray mask to a lesser extent. In contrast, the convergent runs in the pendulum environment showed barely any degradation compared to unsafe training due to the simplicity of the environment.
For boundary projection, the regularisation term alleviated most issues as performance was on par with unsafe training. The observed reduction in non-convergent runs suggests that regularisation improves convergence. However, the fact that non-convergent runs persisted rather than being eliminated indicates that the regularisation coefficient may be too small. The observation that non-convergent runs involve more safeguarding interventions than convergent ones supports this assumption. The impact of regularisation is less pronounced for ray masking, as it improved the convergence speed in the quadrotor environment but not to the level of unsafe training. We attribute the more negligible effect to the fact that regularising the ray mask constantly introduces a gradient component towards the centre. In contrast, regularisation only influences policy updates when actions are unsafe for boundary projection.
Better performance from a passthrough gradient is possible since a robust control invariant state set captures most optimal actions in balancing tasks, which retains the gradient correctness while eliminating the gradient decrease in the mapping direction. Since the unaltered ray mask is almost optimal in the pendulum environment, performance increases are only visible in the quadrotor environment where the gap to unsafe training is closed. The non-convergence of some passthrough runs could be due to the effectively increased learning rate.
The hyperbolic ray mask produces a similar mapping distance to boundary projection due to the hyperbolic tangent function, leading us to expect comparable performance. Unlike boundary projection, the hyperbolic map ensures that a gradient is always available. However, for unsafe actions, this gradient remains small. We observe marginally more stable convergence but significantly lower policy quality than boundary projection. This result is unexpected and may be attributed to the diminished gradient in the mapping direction, as indicated by the frequent safeguarding interventions. The performance of the regularised, hyperbolic ray mask supports this statement, as it achieved the best performance in the quadrotor environment and converged in all ten runs. The large confidence interval and poor mean performance in the pendulum environment were attributed to a single outlier, which was convergent but significantly slower than all other runs.
6.4 COMPARISON OF SAFE CENTRE APPROXIMATIONS
We also compare safe centre approximations for the ray mask. Since the orthogonal approximation only applied to unsafe actions, safe actions were not mapped. In the pendulum task, the one-dimensional action space allows for exact safe centre approximations, which condenses the comparison to rarer interventions by the orthogonal approximation versus the smoother map of the zonotopic approximation. The continuous map the zonotopic approximation provides produced superior final policies and converged faster. The low number of steps of the orthogonal approximation in the quadrotor task was an artefact of the worse policy, as seen in the full figures in Section 7.5. Therefore, smooth safeguards appear more important than rare interventions for learning performance.
Table 5: Comparison of the safe centre approximations.
Pendulum
Modification # Steps Reward # Stuck
Zonotopic 46080 -8.412 0
Orthogonal 71680 -8.491 0
Quadrotor
Modification # Steps Reward # Stuck
Zonotopic 110080 -251.7 0
Orthogonal 5504 -432.7 0
6.5 COMPARISON OF WALL CLOCK TIME
We compare the relative wallclock time of safeguarded training to its unsafe counterpart in Table 6 to estimate the computational overhead. We measured the wall clock time for 10000 steps in the pendulum environment. We observed an at least fourfold increase in computation time with boundary projection. Ray masking took almost double the time of boundary projection, which we trace to the increased computational complexity of its optimisation problems due to the induced safe action set. SHAC produced around a quarter of additional computational overhead, as it must maintain the computational graph for backpropagation. The wall clock time poses a significant downside, although a custom, more efficient implementation could mitigate the effects. Moreover, safeguarding via ray mask is significantly cheaper for explicit safe action sets since the safe centre is provided and does not require costly approximation. However, pre-computing the safe state or action set may not be possible depending on the task, which could further increase the computation needed per training iteration.
Table 6: Relative wall clock time of the different safeguards for 10000 steps in the pendulum environment compared to their unsafe versions.
Policy Optimisation Algorithm
Safeguard SHAC PPO SAC
None 1.000 1.000 1.000
Boundary Projection 5.089 4.483 4.774
Ray Mask 9.857 8.100 7.500
7 LIMITATIONS AND CONCLUSION
This work demonstrates the fundamental applicability and effectiveness of safeguards for analytic gradient-based reinforcement learning, unlocking its usage for safely training agents in simulations before deploying them in safety-critical applications. While we showcased the possibility of achieving performance on par or exceeding unsafe training, success depends on the quality and representation of the safe set.
While we utilised zonotopes, the safeguards presented are not limited to this set representation. The only limitation of the representation is the star-shapedness for the ray mask and the ability to solve the relevant containment problems in an efficient and differentiable manner. The general trade-off in the choice of set representation is achieving a tight approximation of the true safe set versus cheap containment problems. Tighter approximations allow for larger time steps due to the larger safe sets. In contrast, simple representations may allow a closed-form solution and compensate by decreasing the time step size.
A limitation of the zonotope representation is its symmetry. Symmetric safe set representations can unnecessarily limit the size of the safe set. The size limitations are evident in an obstacle avoidance scenario, where the closest distance to an obstacle limits the safe set size direction towards and away from the obstacle.
Moreover, solving the containment problems with CVXPY allows for rapid prototyping but may not offer optimal performance compared to custom solvers and formulations, which could decrease the substantial overhead. In general, safeguarding for the sole sake of efficiency requires either an informative, safe set or an expensive simulation since the sample efficiency gains strongly depend on the quality of the safe set. In contrast, the computational overhead depends only on the representation and dimensionality.
The presented safeguards worked well but are likely not optimal. An interesting idea for future work is deriving a general bijective map inspired by the ray mask [RayMasking] and gauge map [Gauge]. This map would again project actions radially towards an interior point of the safe action set, but the optimal interior point could be different from the geometric centre. Different projection centres could be advantageous in cases where the safe action set is adjacent to the corner of the feasible set, which would shrink the space unevenly. In addition, optimising the trade-off between the mapping distance and the gradient strength could improve convergence properties.
[…]
{IEEEbiography}
[[Uncaptioned image]]T. Walter (Member, IEEE) received the B.Eng. degree in Electrical Engineering and Information Technology from the University of Applied Sciences Munich, Munich, Germany, in 2022, and the M.Sc. degree in Computational Science and Engineering (Honour’s track) from the Technical University of Munich, Munich, Germany, in 2025. He joined the Cyber-Physical Systems Group at Technical University of Munich in 2025. His research interests include modular robotics, machine learning, optimisation, and control theory with manufacturing applications. {IEEEbiography}[[Uncaptioned image]] Hannah Markgraf received a B.Sc. degree in mechanical engineering in 2019 and an M.Sc. in automation and control in 2021, both from RWTH Aachen University. She is working toward a PhD in computer science at the Cyber-Physical Systems Group at the Technical University of Munich. Her research interests include reinforcement learning and optimisation with application to energy management systems. {IEEEbiography}[[Uncaptioned image]] Jonathan Külz received a B.Sc. degree in mechatronics and information technology in 2017 png Karlsruhe Institute of Technology and an M.Sc. degree in robotics, cognition and intelligence in 2021 from Technische Universität München, Munich, Germany, where he is currently working toward the Ph.D. degree in computer science. His research interests include robot morphology optimisation and computational co-design. {IEEEbiography}[[Uncaptioned image]] Matthias Althoff is an associate professor in computer science at the Technical University of Munich, Germany. He received his diploma engineering degree in Mechanical Engineering in 2005, and his PhD degree in Electrical Engineering in 2010, both from the Technical University of Munich, Germany. From 2010 to 2012, he was a postdoctoral researcher at Carnegie Mellon University, Pittsburgh, USA, and from 2012 to 2013, an assistant professor at Technische Universität Ilmenau, Germany. His research interests include formal verification of continuous and hybrid systems, reachability analysis, planning algorithms, nonlinear control, automated vehicles, and power systems.
APPENDIX
7.1 ENVIRONMENT DESCRIPTIONS
All environments share some characteristics. The feasible state sets, feasible action sets, and noise sets are axis-aligned boxes centred at the origin. Additionally, the feasible action set is also of unit length. The transition functions integrate their first-order ordinary differential equation using a semi-implicit Euler scheme.
Pendulum The environment possesses two state variables $s=\begin{bmatrix}\theta&\dot{\theta}\end{bmatrix}^{T}$: the angle $\theta\in\left[-\pi,\pi\right]$ and the angular velocity $\dot{\theta}\in\left[-1000,1000\right]$. The feasible action set is one-dimensional and represents the torque $a$. We obtain a robust control invariant state set using the method in Schäfer et al. [RCI-Sets] to induce a safe action set. The noise set is one-dimensional $w\in\left[-0.1,0.1\right]^{T}$. The time derivative of the state is
$\dot{s}(s,a)=\begin{bmatrix}\dot{\theta}\\ \frac{1.5g\sin{\theta}}{l}+\frac{3ca}{ml^{2}}+w\\ \end{bmatrix}$ (65)
with the gravitational acceleration $g$, the length $l$, the mass $m$, and the torque magnitude $c$. The reward function is
$r(s,a)=-\theta^{2}-\frac{\dot{\theta}^{2}}{10}-\frac{a^{2}}{100}\,.$ (66)
Quadrotor The environment possesses eight state variables $s=\begin{bmatrix}x&y&r&\dot{x}&\dot{y}&\dot{r}&x_{0}&y_{0}\end{bmatrix}^{T}$: two agent coordinates, the roll, the respective velocities, and the initial position. The feasible action set is two-dimensional, representing the thrust $a_{0}$ and the roll angle $a_{1}$. Again, we induce the safe action set using a robust control invariant state set. The noise set is two-dimensional with generator lengths $0.1$. The time derivative of the state is
$\dot{s}=\begin{bmatrix}\dot{x}\\ \dot{y}\\ \dot{r}\\ (a_{0}c_{0}+g)\sin{r}+w_{0}\\ (a_{0}c_{0}+g)\cos{r}-g+w_{1}\\ a_{1}c_{1}pd_{2}-pd_{0}r-pd_{1}\dot{r}\\ 0\\ 0\end{bmatrix}$ (67)
with the torque magnitude $c_{0}$, the roll angle magnitude $c_{1}$, and the PID gains $pd_{0-2}$. The reward is
$\begin{split}r(s,a)&=-2.5\sqrt{(x-x_{0})^{2}+(y-y_{0})^{2}}\\ &-\frac{r+\dot{x}+\dot{y}+\dot{r}}{10}-\frac{(a_{0}c_{0}+g)^{2}}{50}\\ &-\frac{(a_{1}c_{1})^{2}}{100}\,.\end{split}$ (68)
7.2 LEARNING CURVES OF ALL LEARNING ALGORITHMS IN NON-SAFEGUARDED TRAINING
Figure 6: Learning curves of all learning algorithms in non-safeguarded training.
7.3 LEARNING CURVES OF ALL LEARNING ALGORITHMS WITH BOUNDARY PROJECTION
Figure 7: Learning curves of all learning algorithms with boundary projection.
7.4 LEARNING CURVES OF ALL LEARNING ALGORITHMS WITH RAY MASK
Figure 8: Learning curves all learning algorithms with ray mask.
7.5 LEARNING CURVES OF BOTH APPROXIMATIONS
Figure 9: Learning curves of safe centre approximation methods.
7.6 JACOBIAN UNIQUENESS OF BOUNDARY PROJECTION
[…]
$\frac{da_{s}}{da}=-\left(G_{\mathcal{A}_{s}}\right)_{:,\mathcal{I}_{i}}\left(\left(\frac{d\gamma}{da}\right)_{\mathcal{I}_{i},:}^{(p)}+\left(\frac{d\gamma}{da}\right)_{\mathcal{I}_{i},:}^{(h)}\right)\,.$
Using the property that $\text{Null}(A^{T}A)=\text{Null}(A)$, any homogeneous solution $\left(\frac{d\gamma}{da}\right)_{\mathcal{I}_{i},:}^{(h)}$ lies in the null space of $\left(G_{\mathcal{A}_{s}}\right)_{:,\mathcal{I}_{i}}$. Therefore, $\left(G_{\mathcal{A}_{s}}\right)_{:,\mathcal{I}_{i}}\left(\frac{d\gamma}{da}\right)_{\mathcal{I}_{i},:}^{(h)}=\textbf{0}$, and the expression simplifies to
$\frac{da_{s}}{da}=-\left(G_{\mathcal{A}_{s}}\right)_{:,\mathcal{I}_{i}}\left(\frac{d\gamma}{da}\right)_{\mathcal{I}_{i},:}^{(p)}\,,$
which depends only on the unique particular solution provided by the pseudoinverse.