### v2

# Article Category

Abstract
The deployment of autonomous robots in safety-critical applications requires safety guarantees. Provably safe reinforcement learning is an active field of research that aims to provide such guarantees using safeguards. These safeguards should be integrated during training to reduce the sim-to-real gap. While there are several approaches for safeguarding sampling-based reinforcement learning, analytic gradient-based reinforcement learning often achieves superior performance from fewer environment interactions. However, there is no safeguarding approach for this learning paradigm yet. Our work addresses this gap by developing the first effective safeguard for analytic gradient-based reinforcement learning. We analyse existing, differentiable safeguards, adapt them through modified mappings and gradient formulations, and integrate them with a state-of-the-art learning algorithm and a differentiable simulation. Using numerical experiments on three control tasks, we evaluate how different safeguards affect learning. The results demonstrate safeguarded training without compromising performance.
{IEEEkeywords}
Safe reinforcement learning, policy optimisation, differentiable simulation, gradient-based methods, constrained optimisation, first-order analytic gradient-based reinforcement learning
footnotetext: Equal contribution
1 INTRODUCTION
The transfer of physical labour from humans and human-operated machines to robots is a long-standing goal of robotics research. Although robots have been successfully deployed in controlled environments, such as factories, their deployment in human proximity remains challenging [1]. One reason is a lack of safety guarantees to ensure that robots do not harm humans or themselves [2].
A fundamental requirement for safe human-robot interaction is the deployment of controllers with provable safety guarantees. This becomes particularly challenging when using reinforcement learning, which often outperforms classical control methods in uncertain, high-dimensional, and nonlinear environments [3, 4]. To avoid costly and slow real-world training, an agent should preferably train in simulation before deployment to real systems [5, 6]. If the system is safety-critical, applying safeguards already during training is desirable to reduce the sim-to-real gap [7, 8]. Otherwise, the unsafeguarded optimisation may converge to a policy that relies on unsafe states or actions. When the deployment safeguard subsequently restricts the policy, it can become suboptimal or even fail for non-convex objective landscapes, as it has not learned alternative, safe solutions [9]. While crafting reward functions that reliably encode safety requirements could theoretically prevent performance degradation in deployment, this is notoriously difficult without introducing unintended incentives [10, 11]. Moreover, safeguards can aid learning by guiding exploration in challenging solution spaces [12, 13].
In recent years provably safe reinforcement learning has emerged as a research field [14, 15]. Current safeguards are applied in conjunction with reinforcement learning algorithms that rely on the policy gradient theorem to estimate reward landscapes [16, 17, 12, 18, 19, 20]. The advent of differentiable physics simulators [21, 22, 23, 24, 25] allows forgoing this estimation, as a differentiable simulator enables the analytical computation of the reward gradient with respect to actions by differentiating through the dynamics. While these simulators require approximations to remain differentiable, maintaining simulation accuracy is possible. Reinforcement learning algorithms that exploit these gradients promise faster training and better performance [26, 27, 28, 29]. However, existing safeguards for sampling-based reinforcement learning can not be naively applied to these algorithms. Furthermore, there are currently no safeguarding attempts tailored to analytic gradient-based reinforcement learning.
Our work combines state-of-the-art analytic gradient-based reinforcement learning algorithms, differentiable safeguards, and differentiable simulations. As differentiable safeguards, we incorporate a range of provably safe set-based safeguarding methods, whose codomain is a subset of a verified safe action set. This set consists by construction only of safe actions. We formalise desirable safeguarding properties in the context of differentiable optimisation and analyse existing methods with respect to these criteria. Based on this analysis, we propose targeted modifications, such as custom backward passes or adapted maps, that enhance the suitability for analytic gradient-based reinforcement learning. We also extend the applicability of one of the safeguards to state constraints.
We evaluate the provably safe approaches in differentiable simulations of various control problems. We observe sample efficiency and final performance that exceeds or is on par with unsafe training and sampling-based baselines.
In summary, our core contributions are:
  • the first provably safe policy optimisation approach from analytic gradients [^1] Code available at https://github.com/TimWalter/SafeGBPO,
  • an in-depth analysis of some suitable safeguards,
  • adapted backward passes, an adapted mapping and extended applicability of these safeguards,
  • and an evaluation on three control tasks, demonstrating the potential of provably safe reinforcement learning from analytic gradients.
2 RELATED WORK
We provide a literature review on the most relevant research areas: analytic gradient-based reinforcement learning, safeguards, and implicit layers that allow the computation of analytical gradients for optimisation-based safeguards.
2.1 ANALYTIC GRADIENT-BASED REINFORCEMENT LEARNING
Analytic gradient-based reinforcement learning relies on a continuous computational graph from policy actions to rewards, which allows computing the first-order gradient of the reward with respect to the action via backpropagation. Relying on first-order gradient estimators often results in less variance than zeroth-order estimators [30], which are usually obtained using the policy gradient theorem. Less variance leads to faster convergence to local minima of general non-convex smooth objective functions [26, 27]. However, complex or contact-rich environments may lead to optimisation landscapes that are stiff, chaotic or contain discontinuities, which can stifle performance as first-order gradients suffer from empirical bias [30]. Using a smooth surrogate to approximate the underlying noisy reward landscape can alleviate this issue [29]. Moreover, naively backpropagating through time [31] can lead to vanishing or exploding gradients in long trajectories [32].
[…]
2.2 SAFEGUARDING REINFORCEMENT LEARNING
Safeguards are generally categorised according to their safety level [35, 15]. Since we seek guarantees, we limit the discussion to hard constraints. Moreover, safeguards for analytic gradient-based reinforcement learning must define a differentiable map from unsafe to safe actions to allow for backpropagation.
Within this field of research, two common approaches for enforcing safety guarantees are control barrier functions[36, 37, 38, 39, 40] and set-based reachability analysis [41, 42, 17]. Both necessitate some form of environment model in their basic form, but allow incorporating models identified from data [43, 44, 45]. By finding a control barrier function for a given system, forward invariance of a safe state set can be guaranteed [36]. While mainly used for control-affine systems, solutions for non-affine systems that rely on trainable high-order control barrier functions exist [39]. Nevertheless, finding suitable candidates for control barrier functions for complex systems is nontrivial, and uncertainty handling remains challenging [43]. Therefore, we employ set-based reachability analysis, which employs enclosing models of the true environment dynamics to compute all possible system states [17]. Containment of the reachable state set in a safe state set can be guaranteed by adjusting reinforcement learning actions via constrained optimisation. If robust control invariant sets [46] and reachset-conformant system identification are used [45], this approach can be applied efficiently to non-affine systems with uncertainties.
Differentiable maps between unsafe and safe actions are required to combine the safeguarding approaches with analytic gradient-based reinforcement learning, which are only available for continuous action spaces. Krasowski et al. [15] present continuous action projection with safe action sets represented by intervals, where straightforward re-normalisation is employed to map from the feasible action set. Stolz et al. [12] generalise this to more expressive sets with their ray mask method. Tabas et al. [19] derive a differentiable bijection based on Minkowski functionals and apply it to power systems. Chen et al. [18] define differentiable projection layers relying on convex constraints. Gros et al. [20] define the mapping as an optimisation problem to determine the closest safe action. While these approaches are, in principle, differentiable, previous work only utilises them to modify policy gradients. We utilise and modify boundary projection [20] and ray masking [12] in particular to modify policy behaviour in a differentiable setting.
2.3 IMPLICIT LAYERS
Defining the safeguards above can often not be done explicitly in closed form. Instead, they can only be formulated implicitly as a separate optimisation problem. Implicit layers [47, 48] enable an efficient backpropagation through the solution of this separate optimisation problem without unrolling the solver steps. They decouple the forward and backward pass and analytically differentiate via the implicit function theorem [49], which allows for constant training memory. Implicit layers are a potent paradigm that can be utilised for the tuning of controller parameters [50], model identification [51], and safeguarding [18]. Given the complexity of general optimisation problems being NP-hard, it is crucial to approach the implicit formulation with diligence. If a restriction to convex cone programs is possible, solutions can be computed efficiently in polynomial time [52], thereby facilitating a swift forward pass [53][54]. Moreover, this allows the utilisation of CVXPY [55, 56] to formulate the problem, which automatically picks an efficient solver and translates the problem to the desired solver formulation.
3 PRELIMINARIES
We briefly introduce reinforcement learning based on analytical gradients, safe action sets, which serve as the notion of provable safety throughout this work, and the zonotope set representation, which we utilise to describe the safe action sets.
3.1 Analytical gradient-based reinforcement learning
Traditionally, deep reinforcement learning learns an action policy based on scalar rewards without assuming access to a model of the environment dynamics. Prominent algorithms such as REINFORCE [57], Proximal Policy Optimisation [58], or Soft Actor-Critic [59] are based on the policy gradient theorem [60]. This theorem provides an estimator for the gradient of the expected return $J(\theta)$ with respect to the policy parameters $\theta$, given by:
$\frac{\partial}{\partial\theta}J(\theta)=\mathbb{E}_{\pi_{\theta}}\left[\sum_{t=0}^{T}\frac{\partial}{\partial\theta}\log\pi_{\theta}(a_{t}\mid s_{t})Q^{\pi}(s_{t},a_{t})\right]\,,$
where $\pi_{\theta}$ is the parameterised policy, and $Q^{\pi}(s_{t},a_{t})$ denotes the action-value function under policy $\pi$. This gradient estimate can be used to optimise the policy via stochastic gradient ascent.
In contrast, analytical gradient-based reinforcement learning aims to replace this sample-based estimator with a direct gradient computed through a differentiable model of the environment. In such cases, the chain rule can be applied to the full reward computation, yielding an analytical first-order estimate of the policy gradient:
$\frac{\partial}{\partial\theta}J(\theta)=\sum_{t=0}^{T}\left(\frac{\partial r(s_{t},a_{t})}{\partial s_{t}}\frac{\partial s_{t}}{\partial\theta}+\frac{\partial r(s_{t},a_{t})}{\partial a_{t}}\frac{\partial a_{t}}{\partial\theta}\right)$
The term $\frac{\partial s_{t}}{\partial\theta}$ requires backpropagation through time, which can become numerically unstable for long trajectories. This problem motivates the introduction of a regularising critic, resulting in the Short-Horizon Actor-Critic algorithm [29]. For a more detailed review of reinforcement learning methods based on analytical gradients, we refer the interested reader to [29, 34, 25].
3.2 SAFE ACTION SETS
To achieve provable safety, the safety of all traversed states and executed actions must be verifiable. Thus, we introduce a subset of the feasible state set $\mathcal{S}$, the safe state set $\mathcal{S}_{s}\subseteq\mathcal{S}$, containing all states that fulfil all safety specifications. Furthermore, we assume that provable safety is in principle possible, i.e. starting from a safe state, there must exist a sequence of safe actions that ensures the safety of all traversed states
[15, Proposition 1]
$\exists\left(a_{0},a_{1},\dots\right):\mathcal{S}_{i+1}(a_{i},s_{i})\subseteq\mathcal{S}_{s}\quad\forall i\in\mathbb{N}\,\forall s_{0}\in\mathcal{S}_{s}\,,$ (1)
where $\mathcal{S}_{i+1}(a_{i},s_{i})$ denotes the next state set, i.e. the set of reachable states when executing action $a_{i}$ in safe state $s_{i}$. The proposition also implies that at any point in such a sequence, there exists a non-empty safe action set
$\mathcal{A}_{s}(s_{i})=\left\{a_{i}\mid a_{i}\in\mathcal{A},\mathcal{S}_{i+1}(a_{i},s_{i})\subseteq\mathcal{S}_{s}\right\}$ (2)
from which a policy can select actions. In this work, we introduce safeguards $g_{\mathcal{A}_{s}}:\mathcal{X}\mapsto\mathcal{Y}$ with domain $\mathcal{X}\supseteq\mathcal{A}$ and codomain $\mathcal{Y}\subseteq\mathcal{A}_{s}$ that map any feasible, policy-selected action $a_{i}\in\mathcal{A}$ to a safe action $g_{\mathcal{A}_{s}}(a_{i})={a_{s}}_{i}\in\mathcal{A}_{s}$. These safeguards are therefore provably safe by construction.
3.3 ZONOTOPES
We use zonotopes to represent safe sets due to their compact representation and closedness under linear maps and Minkowski sums. Zonotopes are convex, restricted polytopes and are defined as
$\mathcal{Z}=\left\{c+G\beta\,|\,\|{\beta}\|_{\infty}\leq 1\right\}=\left<c,G\right>\,$ (3)
with centre $c\in\mathbb{R}^{d}$, generator matrix $G\in\mathbb{R}^{d\times n}$, and scaling factors $\beta\in{\color[rgb]{0,0,0}[-1,1]}^{n}$. Zonotopes with orthogonal generators and $d=n$ are boxes. We utilise the following properties of zonotopes to formulate our safeguards. The Minkowski sum of two zonotopes $\mathcal{Z}_{1},\mathcal{Z}_{2}\subset\mathbb{R}^{d}$ is [46, Eq. 7a]
$\mathcal{Z}_{1}\oplus\mathcal{Z}_{2}=\left<c_{1}+c_{2},\begin{bmatrix}G_{1}&G_{2}\end{bmatrix}\right>\,.$ (4)
Translating a zonotope is equivalent to translating the centre. Linearly mapping by $M\in\mathbb{R}^{{\color[rgb]{0,0,0}m}\times d}$ yields [46, Eq. 7b]
$M\mathcal{Z}=\left<Mc,MG\right>\,.$ (5)
A support function of a set describes the farthest extent of the set in a given direction. The support function of a zonotope in direction $v\in\mathbb{R}^{d}$ is [61, Lemma 1]
$\rho_{\mathcal{Z}}(v)=v^{T}c+\left\lVert G^{T}v\right\rVert_{1}\,.$ (6)
A point $p\in\mathbb{R}^{d}$ is contained in a zonotope if [62, Eq. 6]
$1\geq\min_{\gamma\in\mathbb{R}^{n}}\|\gamma\|_{\infty}\,\text{s.t.}\,p=c+G\gamma\,.$ (7)
Determining the containment of a zonotope in another zonotope is co-NP complete [62], but a sufficient condition for $\mathcal{Z}_{1}\subseteq\mathcal{Z}_{2}$ is [63, Eq. 15]
1 1$\geq$1 ≥ $\!\min_{\gamma\in\mathbb{R}^{n_{2}},\Gamma\in\mathbb{R}^{n_{2}\times n_{1}}}$ $\left\lVert\begin{bmatrix}\Gamma&\gamma\end{bmatrix}\right\rVert_{\infty}$ (8a)
subject to $G_{1}=G_{2}\Gamma$ (8b)
$c_{2}-c_{1}=G_{2}\gamma\,.$ (8c)
Both containment problems are linear.
4 PROBLEM STATEMENT
Our work considers constrained Markov decision processes $(\mathcal{S},\mathcal{A},P_{f},r,\mathcal{A}_{s})$ of the following elements:
  • a feasible state set $\mathcal{S}\subset\mathbb{R}^{d_{\mathcal{S}}}$,
  • a feasible action set $\mathcal{A}\subset\mathbb{R}^{d}$,
  • a transition distribution $P_{f}(s_{i+1}|s_{i},a_{i})$,
  • a continuously differentiable reward function $r(s_{i},a_{i},s_{i+1})=r_{i}$,
  • and a safe action set $\mathcal{A}_{s}\subseteq\mathcal{A}$.
We seek a safeguarded, stochastic policy that maximises the expected, discounted return over a finite horizon $N$
$\pi^{*}(a|s)=\!\operatorname*{argmax}_{\pi(a|s)}\operatorname*{\mathbb{E}}_{\begin{subarray}{c}a_{i}\sim\pi(a_{i}|s_{i})\\ s_{i+1}\sim P_{f}\end{subarray}}\sum_{i=0}^{N}\delta^{i}\,r(s_{i},{a_{s}}_{i},s_{i+1})$ (9)
with the safe action ${a_{s}}_{i}{\color[rgb]{0,0,0}=g_{\mathcal{A}_{s}}(a_{i})}$, the continuously differentiable safeguard $g_{\mathcal{A}_{s}}\colon\mathcal{A}\to\mathcal{A}_{s}$, and discount factor $\delta\in(0,1{\color[rgb]{0,0,0}]}$.
4.1 ENSURING COMPUTATIONAL TRACTABILITY
We impose additional requirements on the problem to ease the computational burden of safeguarding. This concerns the representation of all sets as closed, convex sets, such as zonotopes.
Safeguarding is computationally cheap when the problem setting explicitly provides such a representation for the safe action set. However, safeguarding might be computationally intractable if the safe action set needs to be constructed from a safe state set, as specified in Equation˜2, even if the safe state set is available as a closed, convex set. In these cases, we assume that the next state set can be derived using only computations adhering to disciplined convex programming [64]. This allows to replace any constraint on a safe action $a_{s_{i}}\in\mathcal{A}_{s}$ by the state constraint $\mathcal{S}_{i+1}(a_{s_{i}},s_{i})\subseteq\mathcal{S}_{s}$.
In practice, the next state set is often enclosed via reachability analysis, leading to a conservative under-approximation of the safe action set in the current state. However, maintaining Equation˜1 requires tight enclosures, which is an active field of research for complex systems [44, 65, 66, 46, 67, 68].
One method we discuss in particular is obtaining safe state zonotopes via robust control invariant sets [46], which guarantee the existence of an invariance-enforcing controller that can keep all future states within the safe set. This is achieved by enclosing the transition distribution at the current state by a linear transition function with a noise zonotope $\mathcal{W}=\left<c_{\mathcal{W}},G_{\mathcal{W}}\right>\subset\mathbb{R}^{d_{\mathcal{S}}}$. Such an enclosure can, for example, be obtained using reachset-conformant identification [44], which bounds the remainder of the Taylor expansion. Such a transition function allows a linear computation of the next state set
$\mathcal{S}_{i+1}(a_{i},s_{i})=Ma_{i}+\left<c+c_{\mathcal{W}},G_{\mathcal{W}}\right>\,,$ (10)
where $c$ is the offset an $M$ the Jacobian of the linearisation.
5 METHOD
Figure˜1 shows the general framework for provably safe, analytic gradient-based reinforcement learning. For any policy output, we apply safeguards that map the unsafe action $a_{i}$ to the safe action ${a_{s}}_{i}$. The safe action is executed in the environment, yielding the next state $s_{i+1}$ and reward $r_{i}$. To train the policy, we calculate the gradient of the reward with respect to the policy output
$\frac{\partial r_{i}}{\partial a_{i}}=\left(\frac{\partial r_{i}}{\partial{a_{s}}_{i}}+\frac{\partial r_{i}}{\partial s_{i+1}}\frac{\partial s_{i+1}}{\partial{a_{s}}_{i}}\right)\frac{\partial{a_{s}}_{i}}{\partial a_{i}}\,.$ (11)
Since the policy output and reward depend on the previous state, full backpropagation requires unrolling the trajectory to determine how all previous policy outputs affect the current reward.
Figure 1: The forward (top) pass of the provably safe policy optimisation from analytic gradients describes the integration of the safeguard in-between the policy and environment. The backward pass (bottom) visualises, how we utilise backpropagation to differentiate through the environment and safeguard to obtain the reward gradient with respect to the policy action. It also highlights how this process in principle requires the unrolling of the previous trajectory.
The following subsections detail the safeguard. First, we formulate generally required and desirable properties in the aforementioned differentiable setting. Then, we introduce the two safeguards used in this work: boundary projection (BP) [20] and ray mask (RM) [12].We structure their introduction by first explaining the principle of the safeguard, then analysing its properties, and finally present our modifications.
5.1 REQUIRED AND DESIRED PROPERTIES
A safeguard for our setting must be provably safe as described in Section˜3.2. To allow for backpropagation, it must also guarantee the existence of a Clark generalised derivative [69] everywhere, which implies that the safeguarding is at least of class $C^{0}$.
Beyond the required, there are additional desired properties. First, the safeguarding should be of class $C^{1}$ and provide full-rank Jacobians $\frac{\partial a_{s}}{\partial a}$ everywhere. A rank-deficient Jacobian can diminish the learning signal by reducing its effective dimensionality and incurring information loss. Second, the number of interventions by the safeguarding should be minimal throughout training and inference. Minimal interventions also serve the last desired property of fast computation. In summary, the safeguard must
  1. P1
    map any action to a safe action,
  2. P2
    be subdifferentiable everywhere, and therefore of class $C^{0}$
and should
  1. P3
    be of class $C^{1}$ and provide full-rank Jacobians everywhere, i.e. be a local diffeomorphism,
  2. P4
    intervene rarely, and
  3. P5
    compute quickly.
In the following, we present two safeguards that offer different trade-offs between the desired properties. We summarise the properties of the safeguards in Table˜1.
Table 1: Properties of the unaltered safeguards.
Boundary Projection Ray Mask
Property˜P1 Safety
Property˜P2 Subdifferentiable
Property˜P3
Class $C^{1}$ almost everywhere almost everywhere
Jacobian rank $d-1$($d$)
Property˜P4 Interventions $\forall a\in\mathcal{A}\setminus\mathcal{A}_{s}$ $\forall a\in\mathcal{A}$
Property˜P5 Computational complexity
For $\mathcal{A}_{s}$ 1 Quadratic Problem 1 Linear Problem
For $\mathcal{S}_{s}$wedge $P_{f,lin}$ 1 Quadratic Problem (1 Conic $\vee$ 1 Quadratic) $\wedge$ 1 Linear Problem
5.2 BOUNDARY PROJECTION
The boundary projection safeguard, which was proposed in [20], maps any action to the closest safe action. By definition, it therefore only affects unsafe actions, which are mapped to the nearest boundary point in the safe action set. In Euclidean space, this corresponds to an orthogonal projection to the boundary of the safe action set. We show an exemplary mapping with boundary projection from an unsafe action $a$ to a safe action $a_{s}$ in Figure˜2.
Figure 2: Boundary projection maps unsafe actions to the boundary of the safe action set by determining the closest safe action.
The safeguard $g_{\mathcal{A}_{s},\text{BP}}(a)$ provides the safe action by solving
$\!\min_{a_{s}}$ $\left\lVert a-a_{s}\right\rVert_{2}^{2}$ (12a)
subject to $a_{s}\in\mathcal{A}_{s}\,.$ (12b)
5.2.1 PROPERTIES
The optimisation problem is always solvable given Equation˜1, and the constraint Equation˜12b satisfies Property˜P1. The implicit function theorem [49] provides the Jacobian of the solution mapping to fulfil Property˜P2.
The distance between the initial and mapped action decreases smoothly with the distance from the unsafe action to the boundary until it is zero for safe actions. However, the mapping location can change abruptly between unsafe actions on different sides of the edges of the safe set. This leads to a jump in the gradient landscape, such that it is only $C^{1}$ almost everywhere. In addition, all unsafe actions along a normal vector to the safe action set are projected to the same safe action on the boundary. Formally, any safe action on the boundary $a_{s,\partial\mathcal{A}_{s}}\in\partial\mathcal{A}_{s}$ is the solution to Equation˜12 for all actions
$a=a_{s,\partial\mathcal{A}_{s}}+t\cdot v\quad\forall t,\,\forall v$ (13)
with the normal, unit vectors $v\in\{v\mid v\cdot(a_{s}-a_{s,\partial\mathcal{A}_{s}})\leq 0,\forall a_{s}\in\mathcal{A}_{s}\}\setminus\{0\}$ and a positive scalar $t$. Therefore, the safeguard cannot propagate gradient components in the mapping direction, such that
$\left(\frac{\partial r}{\partial a_{s}}\frac{\partial a_{s}}{\partial a}\right)^{T}v=0$ (14)
which is especially problematic for gradients parallel to $v$, that occur when the optimal action is safe. In such a case, boundary projection eliminates the gradient, keeping the optimisation stuck indefinitely.
Lemma 1.
Let $\mathcal{A}_{s}$ be the known zonotope $c_{\mathcal{A}_{s}},G_{\mathcal{A}_{s}}\right>$ with generator matrix $G_{\mathcal{A}_{s}}\in\mathbb{R}^{d\times n}$, where ($n\geq d$), such that strong duality holds for Equation˜12. Then, for safe actions, the Jacobian of Equation˜12 is the identity and thus has full rank. The Jacobian is a projection matrix of rank at most $d-1$ for unsafe actions.
We obtain a proof by differentiating through the KKT conditions, which we provide in Section˜7.6. Consequently, boundary projection does not fulfil Property˜P3.
Boundary projection only intervenes for the required unsafe actions and therefore adheres to Property˜P4. If the safe action set is explicitly available as a zonotope, Equation˜7 provides the containment constraint Equation˜12b. For a linearised transition distribution the constraint becomes $\mathcal{S}_{i+1}(a_{i},s_{i})\subseteq\mathcal{S}_{s}$, which is Equation˜8. Both yield quadratic problems, which compute quickly; see Property˜P5.
5.2.2 MODIFICATIONS
To regain a gradient component in the mapping direction and improve Property˜P3, we augment the policy loss function $l_{r}(a_{s},s)$ with a regularisation term [18, Eq. 16]
$l(a,s,a_{s})=l_{r}(a_{s},s)+c_{d}\left\lVert a_{s}-a\right\rVert_{2}^{2}\,.$ (15)
The corresponding gradient
$\frac{\partial l}{\partial a}=\frac{\partial l_{r}}{\partial a}+2c_{d}(a_{s}-a)(\frac{\partial a_{s}}{\partial a}-I)$ (16)
points along the projection direction. The coefficient $c_{d}$ scales the regularisation to remain small relative to the original loss $l_{r}(a_{s},s)$ yet large enough to produce a meaningful gradient component in the mapping direction. If $c_{d}$ is too large, the gradient may become uninformative or vanish for unsafe actions. Vanishing may occur if a hyperbolic tangent function squashes the policy output to adhere to a normalised action space. An excessive $c_{d}$ can yield large upstream gradients, which saturate the hyperbolic tangent. In addition to gradient augmentation, the regularisation encourages the policy to favour safe actions from the start, which is desirable for Property˜P4.
5.3 RAY MASK
Figure 3: The ray mask maps actions towards the safe centre in proportion to the safety domain length , the feasible domain length, and the distance from the action to the safe centre.
The ray mask technique [12] maps every action radially towards the centre of the safe action set $c_{\mathcal{A}_{s}}$, as shown in Figure˜3. Using the unit vector $d_{a}=\frac{a-c_{\mathcal{A}_{s}}}{\left\lVert a-c_{\mathcal{A}_{s}}\right\rVert}$, we define the distances from the safe action set centre to the initial action and the boundaries of the safe and feasible action sets as
$\lambda_{a}$ $=\left\lVert a-c_{\mathcal{A}_{s}}\right\rVert$ (17)
$\lambda_{\mathcal{A}_{s}}$ $=\sup\{\lambda\geq 0:c_{\mathcal{A}_{s}}+\lambda d\in\mathcal{A}_{s}\}$ (18)
$\lambda_{\mathcal{A}}$ $=\sup\{\lambda\geq 0:c_{\mathcal{A}_{s}}+\lambda d\in\mathcal{A}\}\,.$ (19)
We introduce the generalised ray mask as
$g_{\mathcal{A}_{s},\text{RM}}(a)=$ (20)
[…]
where $\varepsilon\ll 1$ ensures numerical stability. The mapping function
$w(\lambda_{a},\lambda_{\mathcal{A}_{s}},\lambda_{\mathcal{A}})\colon(0,\lambda_{\mathcal{A}}]^{2}\times\mathbb{R}_{>0}\mapsto(0,1]$ (21)
is chosen such that $\frac{\partial\omega}{\partial\lambda_{a}}>0$, which ensures a convex mapping.
The linear ray mask introduced in [12, Eq. 6] is obtained by setting $w_{\text{lin}}(\lambda_{a},\lambda_{\mathcal{A}_{s}},\lambda_{\mathcal{A}})=\frac{\lambda_{a}}{\lambda_{\mathcal{A}}}$.
In addition to the constraints introduced in the problem statement, ray masking requires a star-shaped safe action set to ensure that the safe centre and the line segment from the safe boundary point to the safe centre lie within the set. All convex sets, including zonotopes, are star-shaped. While a safe action set induced by a safe state set is not necessarily convex, it is convex for linearised transition distributions.
Lemma 2.
Let $\mathcal{S}_{s}$ be an explicitly available zonotope and a linearised transition distribution yields $\mathcal{S}_{i+1}(a_{i},s_{i})$ according to Equation˜10. Then, $\mathcal{A}_{s}$ according to Equation˜2 is convex.
We provide a proof in Section˜7.8.
5.3.1 COMPUTATION
The distances to the safe action set can be computed by
$\!\max_{\lambda_{\mathcal{A}_{s}}}$ $\lambda_{\mathcal{A}_{s}}$ (22a)
subject to $c_{\mathcal{A}_{s}}+\lambda_{\mathcal{A}_{s}}d_{a}\in\mathcal{A}_{s}\,.$ (22b)
The distances to the feasible action sets can be computed equivalently if they are general zonotopes. For axis-aligned boxes, this can be computed explicitly. For explicitly available safe action zonotopes, the safe centre is defined as the centre of the zonotope. For induced safe action sets the safe centre is not readily available. We present two approaches to approximate it: orthogonal and zonotopic approximation. They are visualised in Figure˜4.
Figure 4: Approximation of the safe centre by expanding a contained zonotope (left) and by piercing the safe action set orthogonal to the boundary and taking the midpoint (right).
The zonotopic approach directly approximates the safe action set by inflating the generator lengths of a zonotope. The under-approximated zonotope $\mathcal{Z}_{\mathcal{A}_{s}}$ is the solution to
$\!\max_{c_{\mathcal{A}_{s}},l_{s}}$ $n]{\prod_{i=1}^{n}l_{s,i}}$ (23a)
subject to $\mathcal{Z}_{\mathcal{A}_{s}}=\left<c_{\mathcal{A}_{s}},G_{\mathcal{A}_{s}}\text{diag}(l_{s})\right>$ (23b)
$\mathcal{Z}_{\mathcal{A}_{s}}\subseteq\mathcal{A}$ (23c)
$\mathcal{S}_{i+1}(\mathcal{Z}_{\mathcal{A}_{s}},s_{i})\subseteq\mathcal{S}_{s}$ (23d)
with $n$ uniformally sampled generator directions $G_{\mathcal{A}_{s}}$. Generally, the number of generators should be in the order of magnitude of the action dimension to provide a good approximation. However, $n$ should also not be too large, since we employ the geometric mean as a computationally cheaper proxy for the volume [12]. The issue arises from the uniformly sampled generator directions, which will likely lead to nearly parallel generators for $n\gg d$. However, computing the volume is only equivalent to the geometric mean for orthogonal generators, since the hypercube volume is the $n$th power of the geometric mean. The geometric mean favours spherical zonotopes for nearly parallel generators over elongated ones, which is not necessarily volume-maximising if the proper safe action set is elongated.
The orthogonal approximation does not directly approximate the safe action set. Instead, it pierces the safe action set orthogonal to the boundary and assumes the midpoint as the safe centre. The orthogonal starting point and direction is determined by Equation˜12, which yields $a_{s,\text{BP}}$ and $d_{\perp}=\frac{a_{s,\text{BP}}-a}{\left\lVert a_{s,\text{BP}}-a\right\rVert}$. Next, we reuse Equation˜22 as
$\!\max_{\lambda_{\perp}}$ $\lambda_{\perp}$ (24a)
subject to $a_{s,\text{BP}}+\lambda_{\perp}d_{\perp}\in\mathcal{A}_{s}\,.$ (24b)
Then, the middle point is the safe centre
$c_{\mathcal{A}_{s}}=a_{s,\text{BP}}+\frac{\lambda_{\perp}}{2}d_{\perp}\,.$ (25)
Since Equation˜12 will only yield a different action for unsafe actions, the orthogonal approximation technique is restricted to those and actions that are already safe remain unmodified.
5.3.2 PROPERTIES
The generalised ray mask fulfils Property˜P1, since its codomain is the safe action set. To illuminate this fact, we remark that the function can be examined in one dimension, the direction along the ray, without loss of generality. The two limit points on the ray are $c_{\mathcal{A}_{s}}$, which maps to $g_{\mathcal{A}_{s},\text{RM}}(c_{\mathcal{A}_{s}})=c_{\mathcal{A}_{s}}\in\mathcal{A}_{s}$, and $c_{\mathcal{A}_{S}}+\lambda_{\mathcal{A}}d_{a}$, which maps to $g_{\mathcal{A}_{s},\text{RM}}(c_{\mathcal{A}_{S}}+\lambda_{\mathcal{A}}d_{a})=c_{\mathcal{A}_{s}}+\lambda_{\mathcal{A}_{s}}d_{a}\in\mathcal{A}_{s}$. Gradients to fulfil Property˜P2 are available from backpropagating through Equation˜20 and the implicit function theorem yields $\frac{\partial c_{\mathcal{A}_{s}}}{\partial a}$ $\frac{\partial\lambda_{\mathcal{A}}}{\partial a}$, and $\frac{\partial\lambda_{\mathcal{A}_{s}}}{\partial a}$.
Regarding smoothness, the ray mask safeguard is of class $C^{1}$ almost everywhere except for the safe set edges, as in boundary projection. The Jacobian of a ray mask is full-rank, adhering mostly to Property˜P3.
Lemma 3.
Let $\mathcal{A}_{s}$ be convex. Then, the Jacobian of any ray mask, as in Equation˜20, has full rank.
We obtain a proof by transforming the action space to a spherical coordinate system, which we provide in Section˜7.9. While the ray mask propagates gradient components in the mapping direction, they are still diminished for the linear mapping. This reduction is most obvious in the scenario, where the feasible and safe action set are spheres with coinciding centres and radii $r_{\mathcal{A}}>r_{\mathcal{A}_{s}}$, and the coordinate system is already spherical and centred. In this scenario, the Jacobian in Equation˜76 reduces to
$\frac{\partial a_{s}}{\partial a}=\begin{bmatrix}{\color[rgb]{0,0,0}\frac{r_{\mathcal{A}_{s}}}{r_{\mathcal{A}}}}&\textbf{0}\\ \textbf{0}&I\end{bmatrix}\,$ (26)
which has a trivial eigenspace, as the Jacobian is diagonal. Consequently, the upstream gradient is only modified in the mapping direction by the factor $\frac{r_{\mathcal{A}_{s}}}{r_{\mathcal{A}}}<1$.
Contrary to the boundary projection safeguard, the ray mask applies to all actions, including safe actions. Moreover, the linear mapping distance decreases only linearly with the distance to the safe centre, as the partial derivative is constant in $\lambda_{a}$
$\frac{\partial\omega_{\text{lin}}}{\partial\lambda_{a}}=\frac{1}{\lambda_{\mathcal{A}}}\,.$ (27)
This means that safe actions far from the centre are also substantially altered, therefore Property˜P4 is not strongly adhered.
In regards to Property˜P5 and computational complexity, the actual application of the ray mask in Equation˜20 is negligible. However, computing the safe boundary Equation˜22 is a linear problem, since Equation˜22b is Equation˜7 or Equation˜8 depending on the availability of the safe action set. The safe centre is given by an explicit safe action set, but for induced ones, the approximations can also be costly. For a linearised transition distribution, the zonotopic approach is a conic problem, while the orthogonal approximation requires the solution of one quadratic problem.
5.3.3 MODIFICATIONS
We propose three possible modifications to the linear ray mask to improve its learning properties. First, we can increase the gradient in the mapping direction with the same regularisation term as in Equation˜15 to compensate for the diminished gradient and nudge towards safety.
Second, we can replace the Jacobian with an identity matrix for faster computation and unimpeded gradient propagation. An identity Jacobian retains the correct gradient directions if the reward-maximising action is safe. However, whether the reward-maximising action is safe is generally unknown and depends on the environment. For unsafe reward-maximising actions, the point of convergence would be the safe boundary point on the line $\overline{c_{\mathcal{A}_{s}}a_{r_{max}}}$, which is no longer optimal.
Finally, we propose the hyperbolic mapping function to limit the perturbation of safe actions. We visually compare both mappings in Figure˜5.
Figure 5: One-dimensional illustration of the linear ray mask safeguarding (RM) and the hyperbolic ray mask safeguarding (HRM). Arrows show how exemplary unsafe actions are mapped to the corresponding safeguarded actions. Shaded regions indicate the distribution of safe actions assuming an initially uniform distribution before safeguarding. The linear ray mask maps $\mathcal{A}$ linearly onto $\mathcal{A}_{s}$, whereas the hyperbolic ray mask maps it exponentially, such that unsafe actions are projected closely to the safe boundary and safe actions are perturbed minimally.
The hyperbolic mapping is defined as
$\omega_{\text{tanh}}(\lambda_{a},\lambda_{\mathcal{A}_{s}},\lambda_{\mathcal{A}})=\frac{\tanh{\frac{\lambda_{a}}{\lambda_{\mathcal{A}_{s}}}}}{\tanh{\frac{\lambda_{\mathcal{A}}}{\lambda_{\mathcal{A}_{s}}}}}\,$ (28)
which maps unsafe actions close to the boundary and safe actions close to themselves, since $\omega_{\text{tanh}}(\lambda_{a}\approx\lambda_{\mathcal{A}_{s}})\approx\lambda_{\mathcal{A}_{s}}$ and $\omega_{\text{tanh}}(\lambda_{a}\ll\lambda_{\mathcal{A}_{s}})\approx\frac{\lambda_{a}}{\lambda_{\mathcal{A}_{s}}}$. It is a valid mapping as $w_{\text{tanh}}(\lambda_{a}=\lambda_{\mathcal{A}})=1$, $w_{\text{tanh}}(\lambda_{a}=0)=0$, and
$\frac{\partial\omega_{\text{tanh}}}{\partial\lambda_{a}}=\frac{1-\tanh^{2}{\frac{\lambda_{a}}{\lambda_{\mathcal{A}_{s}}}}}{\lambda_{\mathcal{A}_{s}}\tanh{\frac{\lambda_{\mathcal{A}}}{\lambda_{\mathcal{A}_{s}}}}}>0\,,$ (29)
since $\lambda_{\mathcal{A}_{s}},\lambda_{\mathcal{A}}>0$ and $\tanh^{2}:\mathbb{R}\mapsto[0,1)$. The hyperbolic map maintains the idea of a radial mapping towards the safe centre, while its mapping behaviour is similar to boundary projection in terms of mapping distance; it nonetheless provides a full Jacobian and a smooth mapping for unsafe and safe actions. Due to its similarity to boundary projection, it also has reduced gradients in the ray direction and benefits from a regularisation term.
6 NUMERICAL EXPERIMENTS
This section tests our two main hypotheses:
  1. H1
    Under safeguarding, analytic gradient-based reinforcement learning achieves higher evaluation returns from fewer environment interactions than sampling-based reinforcement learning.
  2. H2
    Enabling our modified safeguards for analytic gradient-based reinforcement learning during training leads to similar or better policies than unsafe training, given the same number of environment interactions.
The following subsections introduce our experimental setup, discuss the main hypotheses, and provide additional insights.
6.1 SETUP
We conducted all experiments using ten different random seeds. Hyperparameters were tuned exclusively for non-safeguarded training and remained consistent throughout the safeguarded experiments [70]. We assessed the quality of the final policy by calculating the return (Return) achieved over a representative evaluation set. We tracked the number of steps until convergence
(# Steps), defined as reaching within 5% of the final policy. Further, we report the mean and a 95% confidence interval computed using bootstrapping for both the return and number of steps. Lastly, we report the number of runs that did not converge within the maximum allowed number of environment interactions (# Stuck). We excluded non-convergent runs from the return and step calculations to ensure clarity.
In our numerical experiments, we vary all three components of the policy optimisation: learning algorithm, safeguarding, and environment.
6.1.1 LEARNING ALGORITHMS
We choose the first-order reinforcement learning algorithm SHAC [29] over its successor Adaptive-Horizon Actor-Critic [34] due to its maturity, stable convergence, and the lack of stiff dynamics in our tasks. We compare it with two well-established sampling-based reinforcement learning algorithms: on-policy Proximal Policy Optimisation (PPO) [58] and off-policy Soft Actor-Critic (SAC) [59].
Replacing unsafe actions with safe actions inside the policy poses problems for stochastic policies, which rely on the probabilities of the actions. We therefore implement safeguarding as a post-processing step to the policy output without explicitly informing the sampling-based learning processes. This requires them to learn the dynamics associated with the safeguarded environment.
6.1.2 SAFEGUARDS
We evaluate the base versions of the boundary projection and ray mask as safeguards, where we approximate the safe centre using the zonotopic approach. We also assess all modifications to the safeguards individually, as well as the combination of regularisation and the hyperbolic ray mask. We did not investigate all combinations of modifications due to computational constraints.
6.1.3 ENVIRONMENTS
We study three model problems which are detailed in Section˜7.1. The first two are balancing tasks for a pendulum and a quadrotor, where we minimise the distance to an equilibrium position. The safety constraints comprise both action and state constraints that limit, for example, angles and angular velocities. To guarantee constraint satisfaction for all times, we use robust control invariant sets [46] as time-invariant safe state sets. The third environment features an energy management system for a battery and a heat pump aimed at minimising the electricity cost of a household while maintaining a comfortable indoor temperature. Both the state of charge of the battery and the room temperature have limits that must be enforced at all times. When considering the full action range, we achieve this by computing a safe state set that ensures the system can be steered back into the feasible set within one time step.
We build our differentiable simulations according to the gymnasium framework [71] and differentiate through the dynamics using PyTorch’s auto-differentiation engine [72]. We formulate the convex optimisation problems with CVXPY [55, 56] and backpropagate through them with CVXPYLayers [48].
6.2 EVALUATION OF LEARNING ALGORITHMS
In this subsection, we evaluate Hypothesis˜H1 by comparing the sampling-based reinforcement learning algorithms PPO and SAC with the analytic gradient-based algorithm SHAC in safeguarded training.
We first compared the learning algorithms in unsafe training to establish a baseline. We display the key metrics in Table˜2 and the learning curves in Section˜7.2. SHAC converged to the best policies in the pendulum and quadrotor tasks, where it was the only algorithm to balance the quadrotor consistently with minimal effort. However, it performed substantially worse on the energy system task. PPO performed best in energy systems but worst in the pendulum and quadrotor environments.
We attribute the performance degradation of SHAC in the energy system task to the high stochasticity of the environment. Environmental noise likely disrupts the computation of meaningful analytic gradients, and the smooth surrogate critic employed by SHAC may poorly approximate the true reward landscape. In contrast, PPO does not rely on a smooth reward approximation and benefits from the large number of simulation interactions available in the energy system task. Nevertheless, both PPO and SHAC had runs that failed to learn a meaningful policy in the energy system environment. This was also the case for one SHAC run in the pendulum environment.
Table 2: Comparison of learning algorithms in unsafe training.
Environment Algorithm # Step Return # Stuck
Mean 95% CI Mean 95% CI
Pendulum SHAC 12800 [10808, 15360] -8 [-8, -8] 1 / 10
SAC 8513 [6260, 10767] -14 [-15, -11] 0 / 10
PPO 81600 [81600, 81600] -596 [-1174, 300] 0 / 10
Quadrotor SHAC 20364 [11558, 28070] -157 [-169, -140] 0 / 10
SAC 80628 [60096, 109174] -1046 [-1855, 170] 0 / 10
PPO 80640 [42560, 116480] -1710 [-2128, -1267] 0 / 10
Energy System SHAC 259600 [89393, 400400] -114164 [-146048, -79760] 1 / 10
SAC 674999 [554974, 781998] -5225 [-9424, 353] 0 / 10
PPO 748800 [645120, 861120] -2739 [-4598, -167] 2 / 10
After establishing the baselines in unsafe training, we proceed to the hypothesis, comparing safeguarded training. We show the key metrics in Table˜3 and the learning curves in Section˜7.3 and Section˜7.4. On the pendulum task, SAC initially reached near-optimal performance within the first evaluation but later diverged. Under uninformed safeguarding, SAC should benefit from its off-policy nature, but its reliance on the probability of the chosen action outweighs this effect. This issue was most noticeable in the critic loss, which was continuously divergent. The reported lower number of steps for SAC in the pendulum environment is an artefact of the inferior policy. On the quadrotor task, SAC learned ineffectively until the buffer reset roughly twice, at which point a jump in performance was consistently visible. The poor initial performance could result from uninformative earlier samples, although the underlying reason for the drastic performance increase is unclear. PPO benefited from the guided exploration in the pendulum and quadrotor environments, as safety is strongly tied to reward. In the energy system environment, ray masking for PPO significantly hinders learning. This is likely attributed to optimal actions often lying on the boundary of the safe action set in this task. The non-convergent runs of SHAC with boundary projection can be attributed to a total loss of gradient information as outlined in Equation˜14 since the action space in the pendulum environment is one-dimensional. Our observations support the hypothesis since the final policy and convergence speed of SHAC remained superior in the pendulum and quadrotor, while it narrowed the gap in the energy system.
Table 3: Comparison of learning algorithms in safeguarded training.
Environment Safeguard Algorithm # Step Return # Stuck
Mean 95% CI Mean 95% CI
Pendulum BP SHAC 23360 [18560, 28800] -8 [-8, -8] 2 / 10
SAC 2504 [2504, 2504] -1083 [-1103, -1061] 0 / 10
PPO 80240 [78880, 82960] -10 [-10, -9] 0 / 10
RM SHAC 27392 [20992, 33280] -8 [-8, -8] 0 / 10
SAC 2504 [2504, 2504] -424 [-465, -384] 0 / 10
PPO 76160 [72080, 80240] -12 [-12, -11] 0 / 10
Quadrotor BP SHAC 45683 [16498, 74854] -333 [-394, -265] 0 / 10
SAC 110176 [90131, 123196] -338 [-368, -308] 0 / 10
PPO 118720 [89600, 152320] -402 [-453, -350] 0 / 10
RM SHAC 67148 [31909, 99636] -251 [-307, -197] 0 / 10
SAC 80128 [59081, 107171] -377 [-415, -337] 0 / 10
PPO 127680 [107520, 156800] -379 [-419, -330] 0 / 10
Energy System BP SHAC 366960 [213807, 509553] -89167 [-111791, -62775] 0 / 10
SAC 491000 [290000, 685050] -150179 [-252908, -33560] 0 / 10
PPO 661577 [457426, 905307] -2293 [-3421, -906] 3 / 10
RM SHAC 709280 [579920, 840433] -8793 [-12685, -3679] 0 / 10
SAC 355999 [187974, 503073] -1843 [-2396, -1279] 0 / 10
PPO 548352 [313344, 801907] -339006 [-485334, -197690] 0 / 10
6.3 EVALUATION OF SAFEGUARDS
Next, we evaluate Hypothesis˜H2 by comparing the various safeguards introduced in Section˜5 to unsafe training on SHAC. Figure˜6 shows the aggregated learning curves; we report the number of non-convergent runs in Table˜4. Compared to unsafe training, we expect performance to decline for the unaltered safeguards when the optimal action is safe, as is the case for most actions in balancing scenarios. The impact should be more severe for the unaltered boundary projection than for the ray mask, attributed to the lack of gradient propagation in the mapping direction.
Figure 6: Comparison of SHAC in unsafe and safeguarded training via boundary projection (left) and ray mask (right). The modifications are indicated by a different line style. At least one of our safeguards achieves comparable performance to unsafeguarded training.
Table 4: Number of non-convergent runs for the various safeguards.
Safeguard # Stuck
Pen Quad ES
BP Base 2 / 10 0 / 10 0 / 10
Regularised 1 / 10 0 / 10 0 / 10
RM Base 0 / 10 0 / 10 0 / 10
Regularised 0 / 10 0 / 10 0 / 10
Passthrough 2 / 10 1 / 10 2 / 10
HRM Base 1 / 10 0 / 10 0 / 10
Regularised 0 / 10 0 / 10 0 / 10
The impeded learning performance was visible in the quadrotor environment as a drop in policy quality and convergence speed for boundary projection, as the agents were learning very slowly or completely stalled for several individual runs. We made the same observation for the ray mask to a lesser extent. In contrast, the convergent runs in the pendulum environment showed barely any degradation compared to unsafe training due to the simplicity of the environment. Furthermore, ray masking improved performance in the energy system environment substantially, which can be attributed to the guided exploration.
For the boundary projection, the regularisation term alleviated most issues as performance was on par with unsafe training. The observed reduction in non-convergent runs suggests that regularisation improves convergence. However, the fact that non-convergent runs persisted rather than being eliminated indicates that the regularisation coefficient may be too small. The observation that non-convergent runs involve more safeguarding interventions than convergent ones supports this assumption. The impact of regularisation is less pronounced for ray masking, as it improved the convergence speed in the quadrotor environment, but not to the level of unsafe training. We attribute the negligible effect to the fact that regularising the ray mask constantly introduces a gradient component towards the centre. In contrast, regularisation only influences policy updates when actions are unsafe for boundary projection.
Better performance from a passthrough gradient is possible since a robust control invariant state set captures most of the optimal actions in balancing tasks, which retains the gradient correctness while eliminating the gradient decrease in the mapping direction. Since the unaltered ray mask is almost optimal in the pendulum environment, performance increases are only visible in the quadrotor environment, where the gap to unsafe training is closed. The non-convergence of some passthrough runs could be due to the effectively increased learning rate in the balancing tasks. In the energy system task, not all reward-maximising actions are safe, such that the gradients of the passthrough safeguard can be wrong and therefore stall learning.
The hyperbolic ray mask produces a similar mapping distance to boundary projection due to the hyperbolic tangent function, leading us to expect comparable performance. Unlike boundary projection, the hyperbolic map ensures that a gradient is always available. However, for unsafe actions, this gradient remains small. We observe marginally more stable convergence but significantly lower policy quality than boundary projection in the balancing tasks. This result is unexpected and may be attributed to the diminished gradient in the mapping direction, as indicated by the frequent safeguarding interventions. The performance of the regularised, hyperbolic ray mask supports this statement, as it achieved the best performance in the quadrotor environment and converged in all ten runs. The large confidence interval and poor mean performance in the pendulum environment were attributed to a single outlier, which was convergent but significantly slower than all other runs.
6.4 COMPARISON OF SAFE CENTRE APPROXIMATIONS
We also compare safe centre approximations for the ray mask. Since the orthogonal approximation only applied to unsafe actions, safe actions were not mapped. In the pendulum task, the one-dimensional action space allows for exact safe centre approximations, which condenses the comparison to rarer interventions by the orthogonal approximation versus the smoother map of the zonotopic approximation. The continuous map provided by the zonotopic approximation produced superior final policies and converged faster. The low number of steps of the orthogonal approximation in the quadrotor task was an artefact of the worse policy, as seen in the full figures in Section˜7.5. In the energy system, the orthogonal approach converges faster initially but to a worse policy. Therefore, smooth safeguards appear more critical than rare interventions for learning.
Table 5: Comparison of the safe centre approximations.
Environment Approximation # Step Return # Stuck
Mean 95% CI Mean 95% CI
Pendulum Zonotopic 27392 [21241, 32768] -8 [-8, -8] 0 / 10
Orthogonal 30208 [18432, 39424] -8 [-8, -8] 0 / 10
Quadrotor Zonotopic 67148 [30808, 101287] -251 [-306, -194] 0 / 10
Orthogonal 31372 [5476, 57241] -432 [-482, -371] 0 / 10
Energy System Zonotopic 709280 [579898, 829873] -8793 [-12852, -3797] 0 / 10
Orthogonal 109560 [79200, 138632] -79594 [-88277, -70587] 0 / 10
6.5 COMPARISON OF WALL CLOCK TIME
To estimate computational overhead, we compare the relative wall-clock time of safeguarded training with its unsafe counterpart in Table˜6. We measured the wall clock time for 10000 steps in the pendulum environment. We observed at least a four-fold increase in computation time with boundary projection. Ray masking took almost double the time of boundary projection, which we trace to the increased computational complexity of its optimisation problems due to the induced safe action set. SHAC produced around a quarter of the additional computational overhead, as it must maintain the computational graph for backpropagation. The wall clock time poses a significant downside, although a custom, more efficient implementation could mitigate the effects. Moreover, safeguarding via a ray mask is significantly cheaper for explicit safe action sets since the safe centre is provided and does not require costly approximation. However, pre-computing the safe state or action set may not be possible depending on the task, which could further increase the computation needed per training iteration.
Table 6: Relative wall clock time of the different safeguards for 10000 steps in the pendulum environment compared to their unsafe versions.
Learning Algorithm
Safeguard SHAC PPO SAC
No Safeguard 1.000 1.000 1.000
Boundary Projection 5.089 4.483 4.774
Ray Mask 9.857 8.100 7.500
7 LIMITATIONS AND CONCLUSION
This work demonstrates the fundamental applicability and effectiveness of safeguards for analytic gradient-based reinforcement learning, unlocking its usage for safely training agents in simulations before deploying them in safety-critical applications. While we showcased the possibility of achieving performance on par or exceeding unsafe training, success depends on the quality and representation of the safe set.
While we utilised zonotopes, the safeguards presented are not limited to this set representation. The only limitation of the representation is the star-shapedness for the ray mask and the ability to solve the relevant containment problems in an efficient and differentiable manner. The general trade-off in the choice of set representation is achieving a tight approximation of the true safe set versus cheap containment problems. Tighter approximations allow for larger time steps due to the larger safe sets. In contrast, simple representations may allow a closed-form solution and compensate by decreasing the time step size.
A limitation of the zonotope representation is its symmetry. Symmetric safe set representations can unnecessarily limit the size of the safe set. The size limitations are evident in an obstacle avoidance scenario, where the closest distance to an obstacle limits the safe set size direction towards and away from the obstacle.
Moreover, solving the containment problems with CVXPY allows for rapid prototyping but may not offer optimal performance compared to custom solvers and formulations, which could decrease the substantial overhead. The group behind CVXPY recently addressed this issue with a parallel interior point solver [73] and CVXPYgen, which generates a custom solver in C [74], which should achieve substantial speed-ups in future work.
In general, safeguarding for the sole sake of efficiency requires either an informative, safe set or an expensive simulation since the sample efficiency gains strongly depend on the quality of the safe set. In contrast, the computational overhead depends only on the representation and dimensionality.
The presented safeguards worked well but are likely not optimal. An interesting idea for future work is deriving a general bijective map inspired by the ray mask [12] and gauge map [19]. This map would again project actions radially towards an interior point of the safe action set, but the optimal interior point could be different from the geometric centre. Different projection centres could be advantageous in cases where the safe action set is adjacent to the corner of the feasible set, which would shrink the space unevenly. In addition, optimising the trade-off between the mapping distance and the gradient strength could improve convergence properties.
APPENDIX
7.1 ENVIRONMENT DESCRIPTIONS
All environments share some characteristics. The feasible state and action sets are axis-aligned boxes. The feasible action set is also of unit length. The transition distribution is given as a first-order ordinary differential equation with additive, bounded noise, which we integrate using an Euler scheme
$P_{f}(s_{i+1}|s_{i},a_{i})=s_{i}+dt\dot{s}_{i}+\mathcal{W}$ (30)
with the noise zonotope $\mathcal{W}$. For the energy system, we utilise an explicit Euler scheme, whereas for the pendulum and quadrotor, we use a semi-implicit Euler scheme
$P_{f}(s_{i+1}|s_{i},a_{i})=\begin{bmatrix}s_{i}\\ \dot{s}_{i}\end{bmatrix}+dt\begin{bmatrix}\dot{s}_{i+1}\\ \ddot{s}_{i}\end{bmatrix}+\mathcal{W}\,,$ (31)
where we exploit the form of our states.
Pendulum The environment possesses a feasible state set $S=\left[-\pi,\pi\right]\times\left[-8,8\right]$ with the state $s=\begin{bmatrix}\theta&\dot{\theta}\end{bmatrix}^{T}$ representing the angle $\theta$ and the angular velocity $\dot{\theta}$. The feasible action set is $\mathcal{A}=[-1,1]$ with the action $a$ representing the torque. The time derivative of the state is
$\dot{s}(s,a)=\begin{bmatrix}\dot{\theta}\\ \frac{1.5g\sin{\theta}}{l}+\frac{3ca}{ml^{2}}+w\\ \end{bmatrix}$ (32)
with the gravitational acceleration $g$, the length $l$, the mass $m$, and the torque magnitude $c$. The noise zonotope is $\mathcal{W}=\left<\begin{bmatrix}0\\ 0\end{bmatrix},\begin{bmatrix}0&0\\ 0.1&0\end{bmatrix}\right>$. The reward function is
$r(s,a)=-\theta^{2}-\frac{\dot{\theta}^{2}}{10}-\frac{a^{2}}{100}\,.$ (33)
The reward function encodes the goal of balancing the pendulum upright. We define the safety constraints as the part of the state space from which the controller can maintain balance, effectively limiting the velocity and angle close to the upright position. We induce a safe action set from a robust control invariant (RCI) state set, which we obtain by the method in Schäfer et al. [46]. Since the RCI set accounts for errors in the Taylor expansion, we linearise the dynamics but retain the original noise set $\mathcal{W}$ without expansion.
Quadrotor The environment possesses a feasible state set
$\mathcal{S}=[-8,8]^{2}$$\times\left[-\frac{\pi}{12},\frac{\pi}{12}\right]\times[-0.8,0.8]$ (34)
×$\times[-1.0,1.0]\times\left[-\frac{\pi}{2},\frac{\pi}{2}\right]$
[…]
$\dot{s}=\begin{bmatrix}\dot{x}\\ \dot{y}\\ \dot{r}\\ (a_{0}c_{0}+g)\sin{r}+w_{0}\\ (a_{0}c_{0}+g)\cos{r}-g+w_{1}\\ a_{1}c_{1}pd_{2}-pd_{0}r-pd_{1}\dot{r}\end{bmatrix}$ (35)
with the torque magnitude $c_{0}$, the roll angle magnitude $c_{1}$, and the PID gains $pd_{0-2}$. The noise zonotope is $\mathcal{W}=\left<\textbf{0},I\begin{bmatrix}0.1&0.1&0&0&0&0\end{bmatrix}^{T}\right>$, where we denote the diagonalisation of the vector by the multiplication with the identity. The reward is
$\begin{split}r(s,a)&=-2.5\sqrt{(x-x_{0})^{2}+(y-y_{0})^{2}}\\ &-\frac{r+\dot{x}+\dot{y}+\dot{r}}{10}-\frac{(a_{0}c_{0}+g)^{2}}{50}\\ &-\frac{(a_{1}c_{1})^{2}}{100}\,,\end{split}$ (36)
where $x_{0},y_{0}$ encode the position of the quadrotor at reset. The reward function encodes balancing the quadrotor around its spawning location. Similar to the pendulum, the safety constraints prevent escalation of the position and velocity, such that balancing always remains possible within the finite time horizon of the problem. Again, we induce a safe action set from an RCI set using the same approach as the pendulum.
Energy Management System The system has the feasible state set $\mathcal{S}=[0,10]\times[18,24]\times[10,100]$ with the state $s=\begin{bmatrix}e&\vartheta^{\text{in}}&\vartheta^{\text{ret}}\end{bmatrix}$, where $e$ is the charge of the battery, $\vartheta_{\text{in}}$ is the indoor temperature of the building, and $\vartheta_{\text{ret}}$ is the return temperature of the floor heating system. The feasible action set is $\mathcal{A}=[-1,1]^{2}$, representing the power set point for the battery $a_{0}$ and the heat pump $a_{1}$. The time derivative of the state is
$\dot{s}=\begin{bmatrix}a_{0}\\ -c_{0}\vartheta^{\text{in}}+c_{1}\vartheta^{\text{ret}}\\ c_{2}\vartheta^{\text{in}}-c_{2}\vartheta^{\text{ret}}+c_{3}a_{1}\end{bmatrix},$ (37)
where the computation of the coefficients $c_{1-3}$ is detailed in [75].We assume a constant coefficient of performance but model the outdoor temperature and the resulting coefficient as a random variable bounded by the noise zonotope
$\mathcal{W}=\left<\begin{bmatrix}0\\ 10.8986\\ 0\end{bmatrix},\begin{bmatrix}0&0&0\\ 21.1985&0&0\\ 0&0&0\end{bmatrix}\right>\,$ (38)
resulting from the replayed data for the outdoor temperature. The reward is
$r(s,a)=$ $-(a_{0}+a_{1}+p^{\ell}-p^{\text{PV}})dt\,\phi$ (39)
$-100(\vartheta^{\text{in}}-\vartheta^{\text{set}})^{2}\,,$
where $p^{\ell}$ is the inflexible load of the building, $p^{\text{PV}}$ is the output of the photovoltaic generator, $\phi$ is the electricity price, and $\vartheta^{\text{set}}$ is the desired indoor temperature. The reward function encodes the goal of minimising energy expenditure while maintaining room temperature. To facilitate the task the observation $o_{t}=\begin{bmatrix}s_{t}&\vartheta^{\text{out}}_{[t:t+H]}&p^{\text{PV}}_{[t:t+H]}&p^{\ell}_{[t:t+H]}&\phi_{[t:t+H]}\end{bmatrix}$ with the outdoor temperature $\vartheta^{\text{out}}$ additionally contains forecasts besides current measurements. We choose $H=$, resulting in 272727 observations. The safe state set is the subset of the feasible state set from which the agent can recover using the full action range within one time step. Since the state transitions are linear, we do not require an approximation.
7.2 LEARNING CURVES OF ALL LEARNING ALGORITHMS IN NON-SAFEGUARDED TRAINING
Figure 7: Learning curves of SHAC, SAC, and PPO in non-safeguarded training.
7.3 LEARNING CURVES OF ALL LEARNING ALGORITHMS WITH BOUNDARY PROJECTION
Figure 8: Learning curves of SHAC, SAC, and PPO with boundary projection.
7.4 LEARNING CURVES OF ALL LEARNING ALGORITHMS WITH RAY MASK
Figure 9: Learning curves SHAC, SAC, and PPO with ray mask.
7.5 LEARNING CURVES OF BOTH APPROXIMATIONS
Figure 10: Learning curves of SHAC in safeguarded training with the ray mask, where we compare safe centre approximation methods.
7.6 PROOF OF LEMMA 1: JACOBIAN OF BOUNDARY PROJECTION
{IEEEproof}
Equation˜12 becomes
$\min_{a_{s},\gamma}$ $\left\lVert a-a_{s}\right\rVert_{2}^{2}$ (40a)
subject to $a_{s}=c_{\mathcal{A}_{s}}+G_{\mathcal{A}_{s}}\gamma$ (40b)
$\left\lVert\gamma\right\rVert_{\infty}\leq 1$ (40c)
with safe action set $\mathcal{A}_{s}=\left<c_{\mathcal{A}_{s}},G_{\mathcal{A}_{s}}\right>$. Translating Equation˜40 to canonical, quadratic form yields the optimisation variable $z=\begin{bmatrix}a_{s}\\ \gamma\end{bmatrix}\in\mathbb{R}^{d+n}$ and the problem
$\min_{z}$ $\frac{1}{2}z^{T}Qz+q^{T}z$ (41a)
subject to $Az=b$ (41b)
$Kz\leq h$ (41c)
with
$Q$ $=\begin{bmatrix}2I_{d}&\textbf{0}\\ \textbf{0}&\textbf{0}\\ \end{bmatrix}\in\mathbb{R}^{(d+n)\times(d+n)}$ (42)
$q$ $=-2\begin{bmatrix}a\\ \textbf{0}\end{bmatrix}\in\mathbb{R}^{d+n}$ (43)
$A$ $=\begin{bmatrix}I_{d}&-G_{\mathcal{A}_{s}}\end{bmatrix}\in\mathbb{R}^{d\times(d+n)}$ (44)
$b$ $=c_{\mathcal{A}_{s}}\in\mathbb{R}^{d}$ (45)
$K$ $=\begin{bmatrix}\textbf{0}&I_{n}\end{bmatrix}\in\mathbb{R}^{n\times(d+n)}$ (46)
$h$ $=\textbf{1}\in\mathbb{R}^{n}\,,$ (47)
where the subscript of the identity denotes its size, and bold scalars are matrices of appropriate size filled with the scalar. To compute the rank of the Jacobian, we start from the differentials of the KKT conditions [76, Eq. 6]
$\begin{bmatrix}Q&K^{T}&A^{T}\\ \lambda_{D}^{*}K&Kz_{D}^{*}-h_{D}&\textbf{0}\\ A&\textbf{0}&\textbf{0}\end{bmatrix}\begin{bmatrix}\mathop{}\!\mathrm{d}z\\ \mathop{}\!\mathrm{d}\lambda\\ \mathop{}\!\mathrm{d}\nu\end{bmatrix}=$ (48)
$\begin{bmatrix}-\mathop{}\!\mathrm{d}Qz^{*}-\mathop{}\!\mathrm{d}q-\mathop{}\!\mathrm{d}K^{T}\lambda^{*}-\mathop{}\!\mathrm{d}A^{T}\nu^{*}\\ -\lambda_{D}^{*}\mathop{}\!\mathrm{d}Kz^{*}+\lambda_{D}^{*}\mathop{}\!\mathrm{d}h\\ -\mathop{}\!\mathrm{d}Az^{*}+\mathop{}\!\mathrm{d}b\end{bmatrix}$
where the subscript $D$ denotes a diagonalised vector, the superscript * optimal values, $\nu\in\mathbb{R}^{d}$ the dual variables on the equality constraints, $\lambda\in\mathbb{R}^{n}$ the dual variables on the inequality constraints, and $\mathop{}\!\mathrm{d}$ a differential. To obtain the Jacobian with respect to the action, we substitute $\mathop{}\!\mathrm{d}q=\frac{\mathop{}\!\mathrm{d}q}{\mathop{}\!\mathrm{d}a}\mathop{}\!\mathrm{d}a=-2\begin{bmatrix}I_{d}\\ \textbf{0}\end{bmatrix}\mathop{}\!\mathrm{d}a$ and all other differential terms with zero, leaving
$\begin{bmatrix}Q&K^{T}&A^{T}\\ \lambda_{D}^{*}K&Kz_{D}^{*}-h_{D}&\textbf{0}\\ A&\textbf{0}&\textbf{0}\end{bmatrix}\begin{bmatrix}\frac{\mathop{}\!\mathrm{d}z}{\mathop{}\!\mathrm{d}a}\\ \frac{\mathop{}\!\mathrm{d}\lambda}{\mathop{}\!\mathrm{d}a}\\ \frac{\mathop{}\!\mathrm{d}\nu}{\mathop{}\!\mathrm{d}a}\end{bmatrix}=\begin{bmatrix}2\begin{bmatrix}I_{d}\\ \textbf{0}\end{bmatrix}\\ \textbf{0}\\ \textbf{0}\end{bmatrix}\,.$ (49)
We insert the variables and describe the system in expanded form as
$\begin{bmatrix}2I_{d}&\textbf{0}&\textbf{0}&I_{d}\\ \textbf{0}&\textbf{0}&I_{n}&-G_{\mathcal{A}_{s}}^{T}\\ \textbf{0}&\lambda_{D}^{*}&\gamma_{D}^{*}-I_{n}&\textbf{0}\\ I_{d}&-G_{\mathcal{A}_{s}}&\textbf{0}&\textbf{0}\end{bmatrix}\begin{bmatrix}\frac{\mathop{}\!\mathrm{d}a_{s}}{\mathop{}\!\mathrm{d}a}\\ \frac{\mathop{}\!\mathrm{d}\gamma}{\mathop{}\!\mathrm{d}a}\\ \frac{\mathop{}\!\mathrm{d}\lambda}{\mathop{}\!\mathrm{d}a}\\ \frac{\mathop{}\!\mathrm{d}\nu}{\mathop{}\!\mathrm{d}a}\\ \end{bmatrix}=\begin{bmatrix}2I_{d}\\ \textbf{0}\\ \textbf{0}\\ \textbf{0}\\ \end{bmatrix}\,.$ (50)
This yields a coupled system of matrix equations
$2\frac{\mathop{}\!\mathrm{d}a_{s}}{\mathop{}\!\mathrm{d}a}+\frac{\mathop{}\!\mathrm{d}\nu}{\mathop{}\!\mathrm{d}a}$ $=2I_{d}$ (51)
[…]
We solve this system by substitution starting from Equation˜54
$\frac{\mathop{}\!\mathrm{d}a_{s}}{\mathop{}\!\mathrm{d}a}=G_{\mathcal{A}_{s}}\frac{\mathop{}\!\mathrm{d}\gamma}{\mathop{}\!\mathrm{d}a}\,,$ (55)
which we substitute into Equation˜51
$\frac{\mathop{}\!\mathrm{d}\nu}{\mathop{}\!\mathrm{d}a}=2I_{d}-2G_{\mathcal{A}_{s}}\frac{\mathop{}\!\mathrm{d}\gamma}{\mathop{}\!\mathrm{d}a}\,.$ (56)
We substitute this into Equation˜52
$\frac{\mathop{}\!\mathrm{d}\lambda}{\mathop{}\!\mathrm{d}a}=2G_{\mathcal{A}_{s}}^{T}-2G_{\mathcal{A}_{s}}^{T}G_{\mathcal{A}_{s}}\frac{\mathop{}\!\mathrm{d}\gamma}{\mathop{}\!\mathrm{d}a}\,,$ (57)
with which we decouple Equation˜53
$(\lambda_{D}^{*}-2(\gamma_{D}^{*}-I_{n})G_{\mathcal{A}_{s}}^{T}G_{\mathcal{A}_{s}})\frac{\mathop{}\!\mathrm{d}\gamma}{\mathop{}\!\mathrm{d}a}=-2(\gamma_{D}^{*}-I_{n})G_{\mathcal{A}_{s}}^{T}\,.$ (58)
To solve this linear system, we utilise the KKT complementarity slackness conditions in element-wise notation
$\lambda_{i}^{*}(K_{i,:}z_{*}-h_{i})=\lambda_{i}^{*}(\gamma_{i}^{*}-1)=0\quad\forall i=1,\ldots,n\,.$ (59)
The constraint is inactive $\lambda_{i}^{*}=0$ or active $\gamma_{i}^{*}=1$ to fulfil the conditions. We define the index set of active constraints as $\mathcal{I}_{a}=\left\{i\mid\gamma_{i}^{*}=1\right\}$ and of inactive constraints as $\mathcal{I}_{i}=\left\{i\mid\lambda_{i}^{*}=0\right\}$. We examine the rows of Equation˜58
$\lambda_{i}^{*}\left(\frac{\mathop{}\!\mathrm{d}\gamma}{\mathop{}\!\mathrm{d}a}\right)_{i,:}-2(\gamma_{i}^{*}-1)(G_{\mathcal{A}_{s}}^{T}G_{\mathcal{A}_{s}})_{i,:}\frac{\mathop{}\!\mathrm{d}\gamma}{\mathop{}\!\mathrm{d}a}$ (60)
$=-2(\gamma_{i}^{*}-1)(G_{\mathcal{A}_{s}}^{T})_{i,:}\,,$
where we utilise the diagonalised form of the optimal variables and denote full rows or columns with a colon subscript. Next, we combine this with the index sets
$\lambda_{i}^{*}\left(\frac{\mathop{}\!\mathrm{d}\gamma}{\mathop{}\!\mathrm{d}a}\right)_{i,:}$ =0=0= 0 $i\in\mathcal{I}_{a}$ (61)
$(G_{\mathcal{A}_{s}}^{T}G_{\mathcal{A}_{s}})_{i,:}\frac{\mathop{}\!\mathrm{d}\gamma}{\mathop{}\!\mathrm{d}a}$ $=(G_{\mathcal{A}_{s}}^{T})_{i,:}$ $i\in\mathcal{I}_{i}\,.$ (62)
We partition the Jacobian according to the index sets $\frac{\mathop{}\!\mathrm{d}\gamma}{\mathop{}\!\mathrm{d}a}=\begin{bmatrix}\left(\frac{\mathop{}\!\mathrm{d}\gamma}{\mathop{}\!\mathrm{d}a}\right)_{\mathcal{I}_{a},:}\\ \left(\frac{\mathop{}\!\mathrm{d}\gamma}{\mathop{}\!\mathrm{d}a}\right)_{\mathcal{I}_{i},:}\end{bmatrix}$. The assumption of strong duality implies strict complementarity $\lambda_{i}^{*}>0$. Therefore, Equation˜61 specifies the submatrix $\left(\frac{\mathop{}\!\mathrm{d}\gamma}{\mathop{}\!\mathrm{d}a}\right)_{\mathcal{I}_{a},:}=\textbf{0}$. The left-hand side of Equation˜62 reduces to
$(G_{\mathcal{A}_{s}}^{T}G_{\mathcal{A}_{s}})_{i,:}\frac{\mathop{}\!\mathrm{d}\gamma}{\mathop{}\!\mathrm{d}a}$ $=\sum_{j}{(G_{\mathcal{A}_{s}}^{T}G_{\mathcal{A}_{s}})_{i,j}\left(\frac{\mathop{}\!\mathrm{d}\gamma}{\mathop{}\!\mathrm{d}a}\right)_{j,:}}$ (63)
$=\sum_{j\in\mathcal{I}_{i}}{(G_{\mathcal{A}_{s}}^{T}G_{\mathcal{A}_{s}})_{i,j}\left(\frac{\mathop{}\!\mathrm{d}\gamma}{\mathop{}\!\mathrm{d}a}\right)_{j,:}}$
$=\sum_{j\in\mathcal{I}_{i}}{\left(\left(G_{\mathcal{A}_{s}}^{T}\right)_{i,:}\left(G_{\mathcal{A}_{s}}\right)_{:,j}\right)\left(\frac{\mathop{}\!\mathrm{d}\gamma}{\mathop{}\!\mathrm{d}a}\right)_{j,:}}\,,$
using Equation˜61. We isolate the remaining submatrix
$\left(\left(G_{\mathcal{A}_{s}}^{T}\right)_{\mathcal{I}_{i},:}\left(G_{\mathcal{A}_{s}}\right)_{:,\mathcal{I}_{i}}\right)\left(\frac{\mathop{}\!\mathrm{d}\gamma}{\mathop{}\!\mathrm{d}a}\right)_{\mathcal{I}_{i},:}=(G_{\mathcal{A}_{s}}^{T})_{\mathcal{I}_{i},:}\,.$ (64)
The system is always solvable since $n\geq d$, and the right-hand side lies trivially in the column space of the left-hand side. We utilise the Moore-Penrose inverse to solve the system
[…]
$\left(\frac{\mathop{}\!\mathrm{d}\gamma}{\mathop{}\!\mathrm{d}a}\right)_{\mathcal{I}_{i},:}$ $=\left(\left(G_{\mathcal{A}_{s}}^{T}\right)_{\mathcal{I}_{i},:}\left(G_{\mathcal{A}_{s}}\right)_{:,\mathcal{I}_{i}}\right)^{\dagger}\left(G_{\mathcal{A}_{s}}^{T}\right)_{\mathcal{I}_{i},:}\,.$ (66)
We plug this result into Equation˜55
$\frac{da_{s}}{da}$ $=\begin{bmatrix}\left(G_{\mathcal{A}_{s}}\right)_{:,\mathcal{I}_{a}}&\left(G_{\mathcal{A}_{s}}\right)_{:,\mathcal{I}_{i}}\end{bmatrix}\begin{bmatrix}\left(\frac{\mathop{}\!\mathrm{d}\gamma}{\mathop{}\!\mathrm{d}a}\right)_{\mathcal{I}_{a},:}\\ \left(\frac{\mathop{}\!\mathrm{d}\gamma}{\mathop{}\!\mathrm{d}a}\right)_{\mathcal{I}_{i},:}\end{bmatrix}$ (67)
$=\left(G_{\mathcal{A}_{s}}\right)_{:,\mathcal{I}_{i}}\left(\frac{d\gamma}{da}\right)_{\mathcal{I}_{i},:}$
$=\left(G_{\mathcal{A}_{s}}\right)_{:,\mathcal{I}_{i}}\left(\left(G_{\mathcal{A}_{s}}^{T}\right)_{\mathcal{I}_{i},:}\left(G_{\mathcal{A}_{s}}\right)_{:,\mathcal{I}_{i}}\right)^{\dagger}\left(G_{\mathcal{A}_{s}}^{T}\right)_{\mathcal{I}_{i},:}$
The resulting matrix is the projection matrix onto the column space of $\left(G_{\mathcal{A}_{s}}\right)_{:,\mathcal{I}_{i}}$. It is also unique since any homogeneous part of the solution lies in the null space of $\left(G_{\mathcal{A}_{s}}\right)_{:,\mathcal{I}_{i}}$, see Section˜7.7 for details.
The rank of a projection matrix is equal to the design matrix itself $\text{rank}(\frac{\mathop{}\!\mathrm{d}a_{s}}{\mathop{}\!\mathrm{d}a})=\text{rank}(\left(G_{\mathcal{A}_{s}}\right)_{:,\mathcal{I}_{i}})$. Due to the size of $G_{\mathcal{A}_{s}}$, the rank of the Jacobian is at most $d$ when all inequality constraints are inactive, which is only the case for interior points of the zonotope, i.e. safe actions. However, for boundary points, i.e. mapped unsafe actions, at least one constraint is active. Therefore, the rank is at most $d-1$.
Differentiability of Equation˜67 may not be given at points where the active constraint set $\mathcal{I}_{a}$ changes. However, these active set transitions occur on measure-zero sets, and the projection remains of Class $C^{0}$ everywhere. We employ any element from the Clarke subdifferentiable at non-differentiable points, ensuring our approach remains well-defined.
7.7 EXTENSION OF PROOF 1: JACOBIAN UNIQUENESS OF BOUNDARY PROJECTION
[…]
$\frac{da_{s}}{da}=-\left(G_{\mathcal{A}_{s}}\right)_{:,\mathcal{I}_{i}}\left(\left(\frac{d\gamma}{da}\right)_{\mathcal{I}_{i},:}^{(p)}+\left(\frac{d\gamma}{da}\right)_{\mathcal{I}_{i},:}^{(h)}\right)\,.$
Using the property that $\text{Null}(A^{T}A)=\text{Null}(A)$, any homogeneous solution $\left(\frac{d\gamma}{da}\right)_{\mathcal{I}_{i},:}^{(h)}$ lies in the null space of $\left(G_{\mathcal{A}_{s}}\right)_{:,\mathcal{I}_{i}}$. Therefore, $\left(G_{\mathcal{A}_{s}}\right)_{:,\mathcal{I}_{i}}\left(\frac{d\gamma}{da}\right)_{\mathcal{I}_{i},:}^{(h)}=\textbf{0}$, and the expression simplifies to
$\frac{da_{s}}{da}=-\left(G_{\mathcal{A}_{s}}\right)_{:,\mathcal{I}_{i}}\left(\frac{d\gamma}{da}\right)_{\mathcal{I}_{i},:}^{(p)}\,,$
which depends only on the unique particular solution provided by the pseudoinverse.
7.8 PROOF OF LEMMA 2: CONVEXITY OF INDUCED SAFE ACTION SETS
{IEEEproof}
We start by inserting Equation˜10 into Equation˜2
$\mathcal{A}_{s}(s_{i})=\left\{a_{i}\in\mathcal{A}:Ma_{i}+\left<c+c_{\mathcal{W}},G_{\mathcal{W}}\right>\subseteq\mathcal{S}_{s}\right\}\,.$ (68)
Since applying the matrix $M$ is a linear operation and convexity preserving, the safe action set is convex if and only if the set of all translations $\mathcal{T}=\left\{t:t+\mathcal{Z}_{const.}\subseteq\mathcal{S}_{s}\right\}$ is convex. The definition of convexity dictates
$\lambda x+(1-\lambda)y\in\mathcal{S}_{s}$ (69)
for any $x,y\in\mathcal{S}_{s}$ and $\lambda\in\left[0,1\right]$. It also applies to $\mathcal{T}$, since for any $t_{1},t_{2}\in\mathcal{T}$ setting $x=\mathcal{Z}_{const.}+t_{1}$ and $y=\mathcal{Z}_{const.}+t_{2}$ yields
$\lambda(\mathcal{Z}_{const.}+t_{1})+(1-\lambda)(\mathcal{Z}_{const.}+t_{2})$ (70)
$=\mathcal{Z}_{const.}+(\lambda t_{1}+(1-\lambda)t_{2})\in\mathcal{S}_{s}\,,$
which is the respective convexity condition of $\mathcal{T}$.
7.9 PROOF OF LEMMA 3: FULL RANK JACOBIAN OF RAY MASK
{IEEEproof}
Since a ray mask does not modify the action in the centre, the Jacobian for this case is trivially the identity and of full rank. The easiest application and differentiation of the other case is in spherical coordinates, centred at the safe centre. We transform the coordinates with
$a_{o}=\begin{bmatrix}a_{o,r}\\ a_{o,1}\\ \vdots\\ a_{o,n-1}\end{bmatrix}=\text{spherical}(a-c_{\mathcal{A}_{s}})\,,$ (71)
where we adopt the convention of the radius being the first coordinate. In these coordinates, the ray mask modifies a single coordinate
$a_{s,o}=\begin{bmatrix}{\color[rgb]{0,0,0}\omega\frac{\lambda_{\mathcal{A}_{s}}}{\lambda_{a}}}a_{o,r}\\ a_{o,1}\\ \vdots\\ a_{o,n-1}\end{bmatrix}={\color[rgb]{0,0,0}\begin{bmatrix}\omega\lambda_{\mathcal{A}_{s}}\\ a_{o,1}\\ \vdots\\ a_{o,n-1}\end{bmatrix}}\,$ (72)
since $\lambda_{a}=a_{o,r}$ in this coordinate system. Finally, we transform the safe action back
$a_{s}=\text{spherical}(a_{s,o})^{-1}+c_{\mathcal{A}_{s}}\,.$ (73)
The chain rule provides the Jacobian of the ray mask
$\frac{\partial a_{s}}{\partial a}=\frac{\partial a_{s}}{\partial a_{s,o}}\frac{\partial a_{s,o}}{\partial a_{o}}\frac{\partial a_{o}}{\partial a}\,,$ (74)
where the inverse function theorem [49] relates the Jacobians of the transformations as
$\frac{\partial a_{s}}{\partial a_{s,o}}=\frac{\partial a_{o}}{\partial a}^{-1}\,.$ (75)
This relation characterises a similarity transformation, which means $\frac{\partial a_{s}}{\partial a}$ and $\frac{\partial a_{s,o}}{\partial a_{o}}$ are similar and share the same eigenvalues. The Jacobian $\frac{\partial a_{s,o}}{\partial a_{o}}$ is
[…]
which is triangular and, therefore, has the eigenvalues in the diagonal elements
$\sigma_{i}=\begin{cases}{\color[rgb]{0,0,0}\frac{\partial\omega}{\partial a_{o,r}}\lambda_{\mathcal{A}_{s}}}&\text{if }i=1\\ 1&\text{else}\,.\end{cases}$ (77)
Since they are all non-zero and the Jacobian is a square matrix, it has full rank.